### v2

# Utilizing Contextual Clues and Role Correlations for Enhancing Document-level Event Argument Extraction

Abstract
Document-level event argument extraction (EAE) is a vital but challenging task in the domain of information extraction. Most existing approaches focus on the interaction between arguments and event triggers, ignoring two critical points: the information of contextual clues and the semantic correlations among argument roles. In this paper, we propose the CARLG (Contextual Clues Aggregation and Role-based Latent Information Guidance) model, which consists of two modules: Contextual Clues Aggregation (CCA) and Role-based Latent Information Guidance (RLIG), effectively leveraging contextual clues and role correlations for improving document-level EAE. The CCA module adaptively captures and integrates contextual clues by utilizing context attention weights from a pre-trained encoder. The RLIG module captures semantic correlations through role-interactive encoding and provides valuable information guidance with latent role representation. Notably, our CCA and RLIG modules are compact, transplantable and efficient, which introduce no more than 1% new parameters and can be easily equipped on other span-base methods with significant performance boost. Extensive experiments on the RAMS, WikiEvents, and MLEE datasets demonstrate the superiority of the proposed CARLG model. It outperforms previous state-of-the-art approaches by 1.26 F1, 1.22 F1, and 1.98 F1, respectively, while reducing the inference time by 31%. Furthermore, we provide detailed experimental analyses based on the performance gains and illustrate the interpretability of our model.
keywords:
Event Argument Extraction (EAE), Information Extraction (IE), Natural Language Processing (NLP), Natural Language Understanding (NLU).
journal: Knowledge-Based Systems
1 Introduction
Within the domain of Information Extraction (IE), document-level event argument extraction (EAE) is a vital task focused on identifying event arguments and their corresponding roles within textual content associated with specific events. This task is a key process of event extraction (EE) xu2023multi ; lu2022explainable ; lv2022trigger and has broad implications across various fields. For instance, it is instrumental in the analysis of social media networks mitchell1974social ; DING2023110687 ; YANG2022109939 , enhancing recommendation systems  lietal2020gaia ; LI2024111124 ; DONG2022108954 , and improving dialogue systems  Zhang2020.08.03.20167569 ; ZHAO2023110927 . As illustrated in Fig. 1, with the event trigger “homicide” of event type Violence, the goal of EAE system is to identify argument “Freddie Gray” as the role victim, “Baltimore” as the place, and “Officer Caesar Goodson Jr.” as the killer.
Compared with sentence-level EAE liuetal2018jointly ; waddenetal2019entity ; tongetal2020improving , document-level EAE faces more complex issues, notably long-range dependencies and the need for inferences across multiple sentences, discussed in research ebner2020multi and lietal2021document . To solve these problems, some prompt-based methods ma2022prompt ; hsu2023ampere design suitable templates for long text, which solves the long-distance dependency problem and extract arguments in a generative slot-filling manner. Additionally, some works utilize graph structures employing heuristic approaches zhang2020two ; pouranbenveysehetal-2022-document , syntactical frameworks xuetal2021document ; liu2022document ; 9204800 , and Graph Neural Networks kipfsemi for the facilitation of reasoning across sentences. However, current mainstream approaches ignore two essential aspects: (a) the contextual clue information; (b) the semantic correlations among roles.
Contextual clues refer to text elements in the document, excluding the target arguments themselves, which provide beneficial guidance for predicting complex argument roles. These clues play a significant role in event argument extraction by offering valuable information. For instance, in Figure 1, the contextual clues, such as death, accused, and trial, serve as important indicators for identifying the arguments “Freddie Gray” and “Officer Caesar Goodson Jr” within the event Violence. By considering these contextual clues, the model can leverage wider context information to make more informed and accurate predictions. However, many prior studies xu2022two ; lietal2021document ; liu-etal-2023-enhancing-document just rely on the pre-trained transformer encoder to capture global context information in text implicitly, ignoring the necessity of context clues information that is highly relevant to the entity zhou2021document and target event ebner2020multi . Therefore, in this paper, we propose a Contextual Clues Aggregation (CCA) module, which leverages the contextual attention product from the pre-trained transformer encoder to aggregate the event-specific contextual clues for each argument, enhancing the candidate argument representation by incorporating broader and relevant contextual information.
Figure 1: A document from RAMS ebner2020multi dataset. The event violence has the trigger of Homicide and there are three distinct arguments, each assuming different roles. We use the color red to emphasize specific words, representing contextual clues that are significant for the extraction of arguments.
Within the context of real-world document-level EAE task, numerous roles exhibit semantic correlations. For example, as illustrated in Figure 1, there exists a strong semantic correlation between the roles of killer and victim, serving as beneficial guidance for extracting arguments associated with these roles within the target event Violence. In addition, many roles co-occur in multiple events, indicating strong potential for close semantic correlations among them. For a more in-depth analysis, we perform a frequency count and visualize the co-occurrence of the top 15 roles in the RAMS dataset, as shown in Figure 2. Notably, roles such as attacker, target and instrument show high co-occurrence frequency, suggesting their strong semantic correlations compared with other roles. Therefore, in this paper, we propose a Role-based Latent Information Guidance (RLIG) module, comprising role-interactive encoding, the fusion of role information and latent role guidance. Specifically, in role-interactive encoding, we add roles into the input sequence, enabling role embeddings to learn both semantic information and latent correlations among roles through self-attention mechanism. Subsequently, the latent role embeddings are combined with candidate argument spans via pooling and concatenation operations. Finally, we achieve alignment of argument spans, triggers, and roles within a unified space through a novel alignment mechanism utilizing Tucker Decomposition, tucker1964extension , which provides valuable information guidance for document-level EAE.
In summary, to better integrate and leverage our two techniques, we improve existing span-based baselines zhang2020two ; DBLP:journals/corr/abs-1904-05255 for document-level EAE, which has small parameter size and high efficiency. Then the CARLG (Contextual Clues Aggregation and Role-based Latent Information Guidance) model is proposed, by equipping CCA and RLIG modules on our baseline for the aforementioned two crucial points respectively. Significantly, each module leverages the fine-organized attentive weights obtained from the transformer-based pre-trained model, introducing less than 1% of new parameters. Additionally, these two modules can be easily applied to other span-based methods with transformer encoders and provide a significant performance boost in document-level EAE tasks (as shown in our Sec. 5.2), which are compact and transplantable. The computational cost of our model is also lower compared with the recent state-of-the-art approach TSAR, saving no less than 31% of inference time. ur contributions are summarized as follows:
  • 1.
    An enhanced span-based baseline for document-level EAE tasks is introduced, which demonstrates both parameter and computation efficiency, significantly outperforming existing span-based baselines
  • 2.
    We propose a Contextual Clues Aggregation (CCA) module, which adaptively identifies and aggregates the contextual clue information and enhances the representation of candidate arguments by incorporating event-specific contextual information.
  • 3.
    We introduce a Role-based Latent Information Guidance (RLIG) module designed to offer guidance based on latent role information, capturing the semantic correlations among roles.
  • 4.
    Extensive experiments on the public RAMS and WikiEvents lietal2021document datasets demonstrate that the proposed approach combined effectively with various span-based document-level EAE models and lead to significant performance improvements. Our CARLG consistently improves EAE performance on the biological dataset MLEE 10.1093/bioinformatics/bts407 , showing the generality of our approach in different domains.
Figure 2: The visualization represents the co-occurrence frequencies of the top 15 common roles within the RAMS test set, which is defined as the ratio between the co-occurrence times and the sum of the total occurrence times of two roles, and the self-co-occurrence count is set to zero.
2 Related Works
In this section, we provide a brief overview of the related works on the EAE task. We analyze the strengths and weaknesses of current mainstream EAE methods and emphasize the importance of considering contextual clues and role embedding in this task. Additionally, we discuss several applications in the NLP field that leverage contextual clues and relation embedding, which motivates our proposed approach.
2.1 Event Argument Extraction
Event argument extraction (EAE) is pivotal for improving event semantic understanding by identifying event arguments within a context, which constitutes a fundamental step in various downstream applications, including question answering liu2022document ; 6544211 ; 9257099 , text summarization allahyari2017text ; 8949697 , and assisted decision-making zhang2020effect ; 9339934 .
Previous approaches primarily concentrate on identifying the event trigger and its associated arguments from individual sentences. In chenetal2015event , a neural pipeline model was first proposed for event extraction, which was later extended by nguyengrishman2015event ; liuetal2017exploiting to incorporate recurrent neural networks (RNNs) and convolutional neural networks (CNNs) into the pipeline model.  liuetal2018jointly ; yanetal2019event utilize dependency trees to construct semantic and syntactic dependencies and relations, which captures the dependency of words in a sentence. waddenetal2019entity explores all possible spans and utilizes graph neural networks to construct span graphs for information propagation. Additionally, transformer-based pre-trained models  waddenetal2019entity ; tongetal2020improving ; liuetal2022dynamic achieve impressive performance in event extraction.
However, in real-world situations, elements of an event are often expressed across multiple sentences, leading to the emergence of document-level event extraction (DEE) 6910247 ; 7347400 ; SU2024111180 ; SCABORO2023110675 . DEE seeks to identify event arguments throughout whole documents, addressing the issue of extended dependencieswangetal2022query ; xu2022two . We summarize these recent DEE works into 3 categories: (1) Tagging-based methods. Works like wangetal2021cleve ; ducardie2020document apply the BiLSTM-CRF framework for document-level event extraction (DEE). Zheng zhengetal2019doc2edag introduces a mixed graph and a tracking module to capture the interrelations between events. (2) span-based methods. yangetal2021document proposes an encoder-decoder framework for parallel extraction of structured events. Works ebner2020multi ; zhang2020two predict argument roles for selected text segments under a maximum length restriction. Furthermore, Xu xu2022two suggests a dual-stream encoder complemented by an AMR-guided graph to manage extended dependencies. (3) generation-based methods. lietal2021document approaches the issue through conditional generation, whereas duetal2021grit views it as a sequence-to-sequence endeavor. Additionally, Wei weietal2021trigger redefines the task in terms of a reading comprehension challenge. Furthermore, Ren renetal2022clio integrates argument roles into document encoding to account for multiple role details for nested arguments. ma2022prompt ; liu2022document ; XU2023110375 propose a prompt-based method that utilizes slotted prompts for generative slot-filling to extract arguments.
Most existing document-level EAE methods focus on the interaction between arguments and event triggers, ignoring the correlations of roles among events, or the clue information provided by non-arguments context, which can be crucial for the task. Therefore, we attempt to enhance the document-level EAE model with useful contextual information and semantic role correlations.
2.2 Contextual Clues
Contextual clues refer to the information and cues present in the surrounding context of a given entity, event, or situation clark2016contextual ; zhang2020contextual ; liu2022document . These clues provide additional context and can help in understanding the meaning, significance, or relationships associated with the target entity or event peters2018deep . NC-DRE zhang2022nc introduces an encoder-decoder structure, leveraging a pre-trained encoding model and a graph neural network (GNN) for decoding. This setup facilitates the use of non-entity clues in document-level relation extraction. Concurrently, a method for localized context pooling zhou2021document has been put forward, allowing for the direct application of attention mechanisms from pre-trained language models. This method efficiently leverages the contextual information tied to particular entity pairs, aiding in capturing the beneficial context clues for determining relations. Xu xu2021discriminative uses the self-attention mechanism vaswani2017attention to learn a document-level context representation (DLCRep) which contributes to the improvement of information extraction.
Therefore, contextual clues contain important semantic information, which can provide information guidance for deep mining of semantic associations between entities. In this paper, we are the first to incorporate contextual clues techniques into the event argument extraction task and propose the Contextual Clues Aggregation (CCA) module to explicitly capture contextual clues.
2.3 Relation Embedding
Many methods in the field of information extraction leverage relation embeddings to enhance the effectiveness of various tasks 9086164 ; 9694521 ; 9123572 . CRE chen-badlani-2020-relation presents a novel approach to relation extraction by utilizing sentence representations as relation embeddings, which are assessed through entity embeddings trained concurrently with the knowledge base. xuetal2022emrel constructs relation representation explicitly and interacts it with entities and context, which achieves excellent results in multi-triple relation extraction.  han2022documentlevel employs relation embeddings as a medium and introduces two sub-tasks focused on co-occurrence prediction, aiming to enhance the retrieval of relational facts.
Relation embedding is typically utilized in relation extraction tasks. In the case of event argument extraction, we can consider event roles as relations connecting arguments and events. This viewpoint enables the utilization of relation embedding to proficiently capture the semantic correlations between arguments and events within EAE endeavors. Therefore, we aim to use contextual clues and role correlations embedded in latent embedding to strengthen the document-level EAE performance.
3 Enhanced Span-based Baseline
To better integrate our two techniques, we first propose an enhanced span-based baseline for document-level EAE, which exhibits small parameter size and high efficiency. We clarify the preliminaries of the document-level EAE task as follows for better understanding our method.
3.1 Preliminaries
We follow the definition of event argument extraction in RAMSv1.0 benchmark ebner2020multi . To better understand this task, we first clarify some important notions.
  • 1.
    Event mention: a text span within a document that mentions an entity object, e.g., the phrase Paul Duggan in Figure 1.
  • 2.
    Event trigger (trigger): the key term that most explicitly represents the manifestation of the event, e.g., the word homocide in Figure 1.
  • 3.
    Event type: the semantic class of an event, e.g., the event type Violence in Figure 1.
  • 4.
    Argument role (role): the relation between the event and the argument in which it is involved, for example, the role of Baltimore is identified as Place in Figure 1. For this task, argument roles are predetermined for every event type.
  • 5.
    Event argument (argument): an entity mention, value or temporal expression playing a specific argument role, e.g., the argument Freddie Gray in Figure 1.
Figure 3: The architecture or our enhanced baseline.
Formally, for a document $\mathcal{D}$ composed of $N$ words, denoted as $\mathcal{D}=\{w_{1},w_{2},\ldots,w_{N}\}$, a set of pre-defined event types $\mathcal{E}$, a corresponding set of argument roles $\mathcal{R}_{e}$ associated with each event $e\in\mathcal{E}$, and the event trigger $t$ in $\mathcal{D}$. The goal is to detect every pair $(r,s)$ for every event $e$ within the document $\mathcal{D}$, where $r\in\mathcal{R}_{e}$ signifies a role associated with the event $e$, and $s$ denotes a contiguous text span within $\mathcal{D}$. In alignment with earlier research xu2022two ; ebner2020multi , we extract event arguments independently for each event within a document.
3.2 Enhanced Span-based Baseline
In this section, we improve existing span-based baselines ebner2020multi ; zhang2020two and introduce a new baseline for document-level EAE. We integrate other techniques to further improve the performance and the architecture is shown in Figure 3.
3.2.1 Encoder
Following xu2022two ; ebner2020multi , we divide a document into several instances, where each instance only contains one event. For each document instance $\mathcal{D}$, the target event $e$ and the given trigger word $t_{e}$, we wrap the event trigger with special tokens "*""""*"" * " and regard the embedding of the start "*""""*"" * " token as the trigger embedding, which is adapted from the entity marker technique shi2019simple ; soares2019matching . We also add the event type as additional information guidance and the final input sequence $S$ is obtained as follows:
$S\ =\ \mathrm{[CLS]}\ [E_{e}]\ {e}\ [E_{e}]\ \mathrm{[SEP]}\ w_{1}\ w_{2}*\ t_{e}\ *\ ...\ w_{N}\ \mathrm{[SEP]},$
where $[E_{e}]$ is the special event token [^1] In our implementation, using [unused] tokens in BERT devlinetal2019bert and introducing specific tokens for RoBERTa DBLP:journals/corr/abs-1907-11692 . of event $e$. Next, we use a pre-trained transformer encoder to get the embedding of each token as follows:
$\mathbf{H}=\mathrm{Encoder}(\textsc{$S$}).$ (1)
Next, we can get the event representation $\mathbf{H}^{e}\in\mathbb{R}^{1\times d}$ for the start $[E_{e}]$ and the context representation $\mathbf{H}^{w}\in\mathbb{R}^{l_{w}\times d}$, where $l_{w}$ represents the length of the word pieces list. Notably, when the input sequence is longer than the maximum token length, we employ a dynamic window approach zhou2021document ; zhang2021document to process the long sequence. We divide the sequence into overlapping windows and calculate the embeddings for each window. Next, we obtain the final representation by averaging the embeddings of tokens that overlap across various windows.
3.2.2 Candidate Spans Representation
Filtering Impossible Spans   To eliminate noise information from useless spans, we implement a filtering step to reduce the number of candidate spans. Specifically, we exclude argument-impossible spans, such as spans containing commas in the middle. For a document $\mathcal{D}$, candidate spans after argument-impossible exclusion can be derived as $\mathcal{S}=\{s_{i:j}\}$. This enhancement allows us to eliminate approximately 25% of candidate spans on average, enabling our model to focus on candidate spans that contain valuable information.
Candidate Spans Representation  For a candidate argument span ranging from $w_{i}$ to $w_{j}$, we first obtain its average pooling representation $\textbf{s}_{i:j}$ as follows:
$\textbf{s}_{i:j}=\frac{1}{j-i+1}\sum_{k=i}^{j}\mathbf{h}_{k}^{w},$ (2)
where $\mathbf{h}_{k}^{w}$ is the $k^{\mathrm{th}}$ token embedding from $\mathbf{H}^{w}$.
As we extract arguments at the span level, where the boundaries can be ambiguous, we employ fully connected neural networks to construct start and end representations as follows:
$\mathbf{H}^{start}=\mathbf{W}^{start}\mathbf{H},$ (3)
$\mathbf{H}^{end}=\mathbf{W}^{end}\mathbf{H},$ (4)
where $\mathbf{H}$ represents the hidden representation of the input sequence $S$, and $\mathbf{W}^{start},\mathbf{W}^{end}\in\mathbb{R}^{d\times d}$ are trainable parameters. Subsequently, the candidate span representation $\widetilde{\mathbf{s}}_{i:j}$ can be obtained as follows:
$\widetilde{\mathbf{s}}_{i:j}=\mathbf{W}^{s}[\mathbf{h}^{start}_{i};\textbf{s}_{i:j};\mathbf{h}^{end}_{j}],$ (5)
where $\mathbf{W}^{s}\in\mathbb{R}^{3d\times d}$ is the trainable parameter and $\mathbf{h}^{start}_{i}$ and $\mathbf{h}^{end}_{j}$ are the $i^{\mathrm{th}}$ and $j^{\mathrm{th}}$ vector of $\mathbf{H}^{start}$ and $\mathbf{H}^{end}$. we establish the boundary loss for identifying the start and end positions in accordance with xu2022two :
$\begin{array}[]{c}\mathcal{L}_{b}=-\sum_{i=1}^{|\mathcal{D}|}[y_{i}^{s}\log P_{i}^{s}+(1-y_{i}^{s})\log(1-P_{i}^{s})\\ +y_{i}^{e}\log P_{i}^{e}+(1-y_{i}^{e})\log(1-P_{i}^{e})],\end{array}$ (6)
where $y_{i}^{s}$ and $y_{i}^{e}$ denote golden labels and $P_{i}^{s}=\mathrm{sigmoid}(\mathbf{W}_{1}\mathbf{h}^{start}_{i})$ and $P_{i}^{e}=\mathrm{sigmoid}(\mathbf{W}_{2}\mathbf{h}^{end}_{i})$ represent the probabilities of the word $w_{i}$ being predicted as the start or end word of a golden span.
3.2.3 Classification Module
Within each event $e$, for each candidate span $s_{i:j}$, we apply a comparison technique mou-etal-2016-natural ; zeng2020double to enhance its representation. In detail, we concatenate the span representation $\widetilde{\mathbf{s}}_{i:j}$ with the trigger representation $\mathbf{h}_{t}$, their absolute disparity $|\mathbf{h}_{t}-\widetilde{\mathbf{s}}_{i:j}|$, element-wise product $\mathbf{h}_{t}\odot\widetilde{\mathbf{s}}_{i:j}$, the embedding of the event type $\mathbf{H}^{e}$ and the embedding of the span length $\mathbf{E}_{len}$. Finally, through a feed-forward network, we calculate the predicted probability $P(r_{s_{i:j}})$ of the candidate span $s_{i:j}$:
$\mathbf{I}_{i:j}=[\widetilde{\mathbf{s}}_{i:j};\mathbf{h}_{t};|\mathbf{h}_{t}-\widetilde{\mathbf{s}}_{i:j}|;\mathbf{h}_{t}\odot\widetilde{\mathbf{s}}_{i:j};\mathbf{H}^{e};\mathbf{E}_{len}],$ (7)
$P(r_{s_{i:j}})=\mathrm{FFN}(\mathbf{I}_{i:j}).$ (8)
Considering the issue of imbalanced role distribution and the prevalence of negative samples among candidate arguments, we utilize focal loss 8237586 during the training process. Focal loss uses hyperparameters $\alpha$ and $\gamma$ to control the weighting and scaling of the loss function, which allows us to emphasize the importance of positive samples and makes the model focus more on useful information.
$\begin{array}[]{c}\mathcal{L}_{c}=-\sum_{s_{i:j}\in\mathcal{S}}\alpha{{[1-P(r_{s_{i:j}}=y_{s_{i:j}}}))]}^{\gamma}\\ \cdot\log P(r_{s_{i:j}}=y_{s_{i:j}}).\end{array}$ (9)
3.2.4 Training and Inference
During the training process, our model extracts single event at a time. Our training loss consists of two parts: classification loss $\mathcal{L}_{c}$ and boundary loss $\mathcal{L}_{b}$, and $\lambda$ is hyperparameter.
$\mathcal{L}=\mathcal{L}_{c}+\lambda\mathcal{L}_{b}.$ (10)
Consistent with the training process, we predict and choose the argmax role for each candidate span at inference time. If more than one argument span is predicted with the same role, we reserve all spans for this role at inference time.
Figure 4: The main structure of CARLG involves an input sequence, encompassing roles and event types, passing through a transformer-based encoder. This encoder generates output in the form of context representations, role representations, and attention heads. The Contextual Clue Aggregation (CCA) module dynamically consolidates contextual clues, while the Role-based Latent Information Generation (RLIG) module generates latent role embeddings and captures role correlations. Subsequently, the context vector is combined with the latent role vector to create the ultimate span representation. Finally, the classification module leverages this representation to forecast argument roles for all potential spans.
4 CARLG Model
In this section, we introduce Contextual Clues Aggregation (CCA) module which adaptively aggregates contextual clues and Role-based Latent Information Guidance (RLIG) module which effectively captures semantic correlations among roles. These two tecniques are then well equipped on our proposed baseline and our CARLG model are shown in Figure 4.
4.1 Contextual Clues Aggregation
The average pooling representation that previous span-based methods xu2022two ; ebner2020multi ; zhang2020two disregards important contextual information from non-argument words. Although the self-attention mechanism of the pre-trained encoder can capture token-level interactions, it lacks specific global interaction tailored to the event and candidate arguments. To address this limitation, we propose a Contextual Clues Aggregation (CCA) module that selects and combines relevant contextual information for each tuple comprising a candidate span and the event trigger word, denoted as $(s_{i:j},t)$. In this module, as shown in Figure 5, we leverage the attentive heads from the pre-trained transformer encoder, which enables the adoption of attentions from the pre-trained language model instead of training new attention layers.
We employ the token-level attention heads $\textbf{A}^{w}\in\mathbb{R}^{H\times l_{w}\times l_{w}}$ obtained from the final transformer layer of the pre-trained language model. Here, $H$ represents the number of attention heads, and $l_{w}$ signifies the length of word pieces. Utilizing these attention heads, we compute the context attention $\textbf{A}^{C}_{i:j}\in\mathbb{R}^{l_{w}}$ for each candidate span spanning from $w_{i}$ to $w_{j}$, by applying average pooling:
$\textbf{A}^{C}_{i:j}=\frac{1}{H(j-i+1)}\sum_{h=1}^{H}\sum_{m=i}^{j}\mathbf{A}^{w}_{h,m}.$ (11)
Then for span-trigger tuple $(s_{i:j},t)$, we get the contextual clue information $\textbf{c}^{s_{i:j}}\in\mathbb{R}^{d}$ which is crucial for the candidate span, by applying multiplication to the attentions and then normalizing the result:
Figure 5: Contextual Clues Aggregation Module.
$\begin{array}[]{c}\textbf{p}^{c}_{i:j}=softmax(\textbf{A}^{C}_{i:j}\cdot\textbf{A}_{t}^{C}\;)\vspace{1.3ex},\\ \textbf{c}^{s_{i:j}}={\mathbf{H}^{w}}^{T}\ \textbf{p}^{c}_{i:j},\end{array}$ (12)
where $\textbf{A}^{C}_{t}\in\mathbb{R}^{l_{w}}$ is the contextual attention of trigger $t$, $\cdot$ means dot product and $\textbf{p}^{c}_{i:j}\in\mathbb{R}^{l_{w}}$ is the computed contextual attention weight vector. $T$ is the transpose symbol.
We can apply our CCA module to other span-based methods by modifying Eq. 2 and fusing the contextual clue information $\textbf{c}^{s_{i:j}}$ into the average pooling representation as follows:
$\textbf{s}^{c}_{i:j}=tanh(\mathbf{W}^{CCA}[\frac{1}{j-i+1}\sum_{k=i}^{j}\mathbf{h}_{k}^{w};\textbf{c}^{s_{i:j}}]),$ (13)
where $\mathbf{W}^{CCA}\in\mathbb{R}^{2d\times d}$ is learnable parameter. $\textbf{s}^{c}_{i:j}$ is the enhanced representation of the candidate span $s_{i:j}$.
Figure 6: Role-based Latent Information Guidance Module. Sub-figure (a) illustrates the information fusion and (b) illustrates the Latent Role Information Guidance operation.
4.2 Role-based Latent Information Guidance
In this section, we propose Role-based Latent Information Guidance (RLIG) module, which is shown in Figure 6.
4.2.1 Role-interactive Encoding
To capture the semantic correlations between roles effectively, we integrate the information of role types into the input and utilize multi-head attention to facilitate interaction between the context and roles. This enables us to obtain the context and role representation in a shared knowledge space. We generate latent role embeddings by assigning unique special tokens   [^2] In our implementation, using [unused] tokens in BERT devlinetal2019bert and introducing specific tokens for RoBERTa DBLP:journals/corr/abs-1907-11692 to embody different roles in the framework of a pre-trained model, with each role type gaining a unique latent representation. Considering that role names carry valuable semantic information wangetal2022query , we encase role names in special tokens indicative of their role types, selecting the embedding of the initial special token as the role’s representation. This approach ensures the capture of unique attributes and subtle semantic distinctions of each role. For instance, the role “Victim” is denoted as $[R_{0}]\quad Victim\quad[R_{0}]$, with $[R_{0}]$ representing the specialized token for “Victim”.
Given the input document $\mathcal{D}$, the target event $e$ and its associated set of roles $\mathcal{R}_{e}=\{{r}_{1},{r}_{2},{r}_{3},...\}$, we enhance the input sequence described in Sec. 3.2.1 by concatenating these elements in the following manner:
$S\ =\ \mathrm{[CLS]}\ [E_{e}]\ {e}\ [E_{e}]\ \mathrm{[SEP]}\ w_{1}\ w_{2}\ *\ t_{e}\ *\ ...\ w_{N}\ \mathrm{[SEP]}$
$[R_{0}]\ {r_{0}}\ [R_{0}]\ [R_{1}]\ {r_{1}}\ [R_{1}]\ ...\ \mathrm{[SEP]}$,
where $[R_{0}]$ and $[R_{1}]$ denote the special role type tokens for $r_{0}$ and $r_{1}$, respectively, and we utilize the final $\mathrm{[SEP]}$ token to signify the ”none” category. By employing role-interactive encoding, we are able to get the event representation $\mathbf{H}^{e}\in\mathbb{R}^{1\times d}$ for the start event $[E_{e}]$, the context representation $\mathbf{H}^{w}\in\mathbb{R}^{l_{w}\times d}$, and the role representation $\mathbf{H}^{r}\in\mathbb{R}^{l_{r}\times d}$, all from $\mathbf{H}^{s}$. Here, $l_{w}$ denotes the length of the list of word pieces, and $l_{r}$ represents the length of the list of roles.
4.2.2 Role Information Fusion
Through role-interactive encoding, the role embeddings are able to capture semantic correlations and dynamically adapt to the target event and context. To ensure each candidate argument benefits from the guidance of role-specific information, we adapt our method for aggregating contextual clues to selectively incorporate role information. The aggregated role information $\textbf{r}^{s_{i:j}}\in\mathbb{R}^{d}$ for ${s_{i:j}}$ is obtained through contextual pooling, which involves adjusting the procedure described in Equations 11 and 12:
$\begin{array}[]{c}\textbf{A}^{R}_{i:j}=\frac{1}{H(j-i+1)}\sum_{h=1}^{H}\sum_{m=i}^{j}\mathbf{A}^{r}_{h,m}\vspace{1.3ex},\\ \textbf{p}^{r}_{i:j}=softmax(\textbf{A}^{R}_{i:j}\cdot\textbf{A}_{t}^{R}\;)\vspace{1.3ex},\\ \textbf{r}^{s_{i:j}}={\mathbf{H}^{r}}^{T}\ \textbf{p}^{r}_{i:j},\end{array}$ (14)
where $\textbf{A}^{r}\in\mathbb{R}^{H\times l_{w}\times l_{r}}$ represents the role attention heads obtained from the final transformer layer of the pre-trained model. $\textbf{A}^{R}_{i:j}\in\mathbb{R}^{l_{r}}$ represents the role attention for each candidate span, while $\textbf{A}^{R}_{t}\in\mathbb{R}^{l_{r}}$ denotes the role attention associated with the trigger $t$. Additionally, the vector $\textbf{p}^{r}_{i:j}\in\mathbb{R}^{l_{r}}$ characterizes the computed attention weight vector for roles. We can modify Eq. 13 and integrate the latent role information $\textbf{r}^{s_{i:j}}$ into the average pooling representation as follows:
$\textbf{s}^{r}_{i:j}=tanh(\mathbf{W}^{RLIG}[\frac{1}{j-i+1}\sum_{k=i}^{j}\mathbf{h}_{k}^{w};\textbf{r}^{s_{i:j}}]),$ (15)
where $\mathbf{W}^{RLIG}\in\mathbb{R}^{2d\times d}$ is learnable parameter.
4.2.3 Latent Role Information Guidance
We first map the original $d$-dimensional role features into a reduced dimension $d^{\prime}$ via a fully connected network to obtain the latent role information $\mathbf{H}^{r^{\prime}}\in\mathbb{R}^{l_{r}\times d^{{}^{\prime}}}$ as follows:
$\mathbf{H}^{r^{\prime}}=\mathbf{H}^{r}\mathbf{W}_{3}+\mathbf{b}_{w},$ (16)
where $\mathbf{W}_{3}\in\mathbb{R}^{d\times d^{{}^{\prime}}}$ and $\mathbf{b}_{w}\in\mathbb{R}^{d^{{}^{\prime}}}$ are learnable parameters. To fully utilize the expressive power of role embeddings, we propose a factorization-based alignment approach using Tucker decomposition  tucker1964extension . We define candidate span-trigger-role triples set $\mathcal{T}=\{<s_{i:j},r_{k},t_{e}>|s_{i:j}\in\mathcal{S},r_{k}\in\mathcal{R}_{e}\}$. Then we introduce a core tensor $\mathbf{Z}\in\mathbb{R}^{d_{i}\times d^{{}^{\prime}}}$ where $d_{i}$ is the hidden dimension of the combined representation $\mathbf{I}_{i:j}$ calculated in Eq. 7, and for simplity, the validity for each triple $<s_{i:j},r_{k},t_{e}>\in\mathcal{T}$ is scored as follows:
$\varphi(s_{i:j},r_{k},t_{e})\;=\;\sigma(\mathbf{Z}\;\times_{1}\;\mathbf{I}_{i:j}\times_{2}\;{\widehat{\mathbf{r}}}_{\mathrm{k}}\;+b_{k}),$ (17)
where ${\widehat{\mathbf{r}}}_{\mathrm{k}}=\mathbf{H}^{r^{\prime}}_{k,:}$ represents the transformed role features and ×$\times_{n}$ denotes the tensor product operation along the $n^{th}$ dimension. The calculation of $\varphi$ for all triples is performed concurrently via a batched tensor product technique, and the model is trained employing focal loss as described in reference 8237586 .
5 Experiments
5.1 Experimental Setup
5.1.1 Datasets
We evaluate the proposed method on three large-scale document-level EAE datasets, that is, RAMS ebner2020multi , WikiEvents lietal2021document and MLEE 10.1093/bioinformatics/bts407 , and detailed data statistics are listed in Table 1.
Table 1: Detailed data statistics.
Dataset RAMS WikiEvents MLEE
# Event types 139 50 23
# Args per event 2.33 1.40 1.29
# Events per text 1.25 1.78 3.32
# Role types 65 57 8
# Events
 Train 7329 3241 4442
 Dev 924 345 -
 Test 871 365 2200
  • 1.
    RAMS  The dataset includes 9,124 annotated events sourced from English online news articles. In this dataset, each event forms a single instance, and events that occur within the same context are grouped together into a single instance containing multiple events. We adopt the original train/dev/test split following ebner2020multi .
  • 2.
    WikiEvents  This dataset centers on events documented in English Wikipedia, encompassing both the events and the related news articles referencing these events. It’s annotated with co-reference links for arguments. However, for our experiments, we only employ the precise argument annotations provided.
  • 3.
    MLEE  This dataset is a document-level event extraction dataset that focuses on bio-medical publications, consisting of manually annotated abstracts written in English. We follow the procedure outlined in 10.1093/bioinformatics/btaa540 and it’s worth noting that the preprocessed dataset only provides a train/test data split, without a dedicated development set. In this case, we use the training set as the development set for our experiments following  10.1093/bioinformatics/btaa540 .
5.1.2 Evaluation Metrics
For the RAMS dataset, we adopt the method detailed in xu2022two , focusing on both Span F1 and Head F1 scores across the development and test sets. Span F1 evaluates the precision of matching predicted argument spans against the gold standard, while Head F1 quantifies the accuracy of identifying the central word in each argument span. For a comprehensive assessment, we also incorporate Span F1 for argument identification (Arg IF) on the test set, as suggested by ma-etal-2022-prompt . Regarding the the WikiEvents dataset, we present both Head F1 and Coref F1 metrics on the test set for Argument Identification (Arg IF) and Argument Classification (Arg CF) tasks, consistent with the methodology described in lietal2021document . The Coref F1 score, inspired by the evaluation technique from jigrishman2008refining , gauges the coreference accuracy between extracted arguments and the established gold standard. For the MLEE dataset, we report Span F1 scores on the test set for both Arg IF and Arg CF tasks, following the methodologies in lietal2021document ; ma-etal-2022-prompt .
5.1.3 Baselines
[…]
Table 2: Detailed hyperparameter settings.
Hyperparameters RAMS WikiEvents MLEE
Training Epochs 50 100 100
Warmup Ratio 0.2 0.05 0.1
Transformer LR 3e-5 3e-5 3e-5
Not Transformer LR 1e-4 1e-4 1e-4
Dropout Ratio 0.1 0.1 0.1
Batch Size 4 4 4
Max Span Length 8 8 8
Max Input Length 1024 1024 1024
Lambda 0.05 0.1 0.05
Alpha 10 10 10
Gamma 2 2 2
5.1.4 Hyperparameter Setting
Due to the fact that our proposed module does not rely heavily on hyperparameters, most of the hyperparameters follow the same configuration of TSAR xu2022two . More detailed hyperparameters settings are listed in Table 2.
Table 3: The main results for the RAMS dataset. Arg IF denotes the F1 score for the Argument Identification subtask, and other metrics are for the Argument Classification subtask. * indicates results reported by  ma-etal-2022-prompt . The highest scores are highlighted in bold, while the second-highest scores are underlined.
Method Dev Test
Span F1 Head F1 Span F1 Head F1 Arg IF
Tag-based
BERT-CRF ebner2020multi 38.1 45.7 39.3 47.1 -
[…]
Generation-based
FEAE weietal2021trigger - - 47.40 - -
[…]
EACE zhou2024eace - - 50.1 56.3
RKDE hu2023role - - 50.3 55.1
Span-based
Two-Step zhang2020two 38.9 46.4 40.1 47.7 -
[…]
${\text{TSAR}}_{{\text{base}}}$ xu2022two 45.23 51.70 48.06 55.04 -
${\text{TSAR}}_{{\text{large}}}$ xu2022two 49.23 56.76 51.18 58.53 -
${\text{SCPRG}}_{{\text{base}}}$ liu-etal-2023-enhancing-document 46.06 52.75 48.18 55.83 -
${\text{SCPRG}}_{{\text{large}}}$ liu-etal-2023-enhancing-document 50.23 57.22 51.88 59.06 -
Our Methods
[…]
${\text{CARLG}}_{{\text{base}}}$ 46.83 53.42 49.81 56.21 55.91
${\text{CARLG}}_{{\text{large}}}$ 50.64 57.71 52.82 59.79 56.71
Span-based+Ours
Two-Step+Ours 43.42 50.58 45.27 52.79 50.34
${\text{TSAR}}_{{\text{base}}}$+Ours 46.52 53.28 49.23 56.07 53.22
${\text{TSAR}}_{{\text{large}}}$+Ours 50.61 57.96 52.52 59.77 57.03
5.2 Main Results
5.2.1 Performance of our enhanced baseline
[…]
5.2.2 Performance of CARLG
[…]
Table 4: Main results of WikiEvents and MLEE datasets. * means we rerun their code based on their experimental settings (hyperparameter tuning is conducted when necessary). The highest scores are highlighted in bold, while the second-highest scores are underlined.
WikiEvents MLEE
Method Arg IF Arg CF Arg IF Arg CF
Head F1 Coref F1 Head F1 Coref F1 Span F1 Span F1
Tag-based
BERT-CRF ebner2020multi 69.83 72.24 54.48 56.72 - -
Generation-based
BERT-QA du-cardie-2020-event 61.05 64.59 56.16 59.36 - -
BERT-QA-Doc du-cardie-2020-event 39.15 51.25 34.77 45.96 - -
EEQA du-cardie-2020-event - - 56.9 - 68.4 66.7
[…]
${\text{EA}^{2}\text{E}}_{{\text{large}}}$ zengetal2022ea2e 74.62 75.77 68.61 69.70 - -
EACE zhou2024eace - 71.1 - 66.2 - -
RKDE hu2023role - 69.1 - 63.8 - -
Span-based
Two-Step zhang2020two 70.32* 72.27* 55.11* 57.12* 70.64* 68.74*
${\text{TSAR}}_{{\text{base}}}$ xu2022two 75.52 73.17 68.11 66.31 72.21* 71.03*
${\text{TSAR}}_{{\text{large}}}$ xu2022two 76.62 75.52 69.70 68.79 73.92* 72.71*
${\text{SCPRG}}_{{\text{large}}}$ liu-etal-2023-enhancing-document 77.52 76.01 70.11 69.25 73.88* 73.21*
Our Methods
[…]
${\text{CARLG}}_{{\text{base}}}$ 76.22 74.97 69.03 68.53 73.26 71.23
${\text{CARLG}}_{{\text{large}}}$ 77.22 76.03 70.92 70.11 75.90 73.71
Span-based+Ours
Two-Step+Ours 73.20 72.96 60.90 61.04 72.42 70.11
${\text{TSAR}}_{{\text{base}}}$+Ours 76.43 75.10 68.81 68.22 73.13 71.12
${\text{TSAR}}_{{\text{large}}}$+Ours 77.46 76.10 70.88 70.28 74.98 73.48
5.2.3 Performance of Span-based Methods equipped with CARLG
In order to evaluate the compatibility of our CCA and RLIG modules, we equip these two modules on some classical span-based methods and conduct experiments on three datasets. As is shown in Table 3 and Table 4, both Two-Step and TSAR exhibit improved performance when equipped with CCA and RLIG modules. Two-Step yields an improvement of +5.17/+5.09 Span F1/Head F1 for Arg CF on RAMS, +2.88/+5.79 Head F1 for Arg IF/Arg CF on WikiEvents, and +1.78/+1.37 Span F1 for Arg IF/Arg CF on MLEE. ${\text{TSAR}}_{{\text{large}}}$ achieves notable improvements in Span F1/Head F1 for Arg CF on RAMS (+1.34/+1.24), Head F1 for Arg IF/Arg CF on WikiEvents (+0.84/+1.18), and Span F1 for Arg IF/Arg CF on MLEE (+1.06/+0.77). It is worth noting that the improvement of TSAR after incorporating CCA and RLIG modules is relatively modest. This could be attributed to the fact that the AMR graph already contains contextual semantic information, which may mitigate the need for additional contextual clue integration. These results illustrate that our two modules can be easily applied to other span-based methods with transformer encoders and provide a significant performance boost in document-level EAE tasks, which are compact and transplantable.
[…]
Method Dev Test
Span F1 Head F1 Span F1 Head F1
[…]
 -Boundary Loss 42.18 50.28 42.33 51.68
 -Focal Loss 44.97 51.36 45.83 53.03
 -FIS 43.27 50.64 45.26 52.11
Table 6: Ablation Study on RAMS and WikiEvents for CARLG. Span, Head and Coref refer to Span F1, Head F1 and Coref F1 scores respectively.
RAMS WikiEvents
Method Params Dev Test Arg IF Arg CF
Span Head Span Head Head Coref Head Coref
${\text{CARLG}}_{{\text{base}}}$ 120.87M 46.83 53.42 49.81 56.21 76.22 74.97 69.03 68.53
-CCA 120.28M 45.78 52.67 47.43 54.57 74.57 73.86 67.44 66.87
-RLIG 118.78M 45.27 52.36 47.91 54.52 75.59 73.83 68.26 67.49
-both 118.19M 45.07 51.59 45.76 53.16 73.92 73.22 66.90 65.98
${\text{CARLG}}_{{\text{large}}}$ 374.70M 50.64 57.71 52.82 59.71 77.22 76.03 70.92 70.11
-CCA 373.65M 49.68 56.46 50.84 57.9 75.68 73.82 69.79 68.65
 -RLIG 371.68M 49.96 57.32 51.28 58.46 76.30 74.08 69.87 68.96
-both 370.63M 48.33 54.04 47.52 55.61 75.45 73.55 68.63 67.30
5.3 Ablation Study
In order to effectively showcase the effectiveness of our components, we carry out an ablation study on CARLG using the RAMS and WikiEvents datasets, as depicted in Table 6. First, when we remove the Contextual Clues Aggregation (CCA) module, we observe a decrease in Span F1 and Head F1 scores of ${\text{CARLG}}_{{\text{base}}}$/ ${\text{CARLG}}_{{\text{large}}}$ by 2.83/1.98 and 1.64/1.81 on the test set, respectively, which indicates that our CCA module plays a crucial role in capturing essential contextual clues for document-level EAE.
Moreover, when removing the Role-based Latent Information Guidance (RLIG) module [^3] We additionally remove the respective role tokens that were included in the input sequence., we observe a significant decrease in performance for ${\text{CARLG}}_{{\text{base}}}$/ ${\text{CARLG}}_{{\text{large}}}$, with a drop of 1.9/1.54 in Span F1 and 1.69/1.07 in Head F1 on the RAMS test set. This demonstrates the effect of the RLIG module in capturing meaningful semantic correlations among roles, improving the performance of document-level EAE. Moreover, when both the CCA and RLIG modules are removed, the decline in performance is even more noticeable in comparison to excluding a single module. This illustrates the synergistic enhancement achieved by the collaboration of these two modules.
[…]
Figure 7: Context weights for a sample from RAMS are displayed, focusing on the span-trigger pair (Islamic State, wounded). Various color intensities are used to depict the attention weights.
5.4 Analysis of Contextual Attentive Weights
To evaluate the validity of our Contextual Clues Aggregation (CCA) module in capturing beneficial contextual clues for candidate argument spans, we present a visualization of the contextual weights $\textbf{p}^{c}_{i:j}$, as detailed in Eq.12, using an instance from the RAMS dataset. As shown in Figure 7, the CCA module assigns high weights to non-argument contextual clues such as attack, responsibility, and terrorist attack, which are highly relevant to the span-trigger pair (Islamic State, wounded). Notably, our Contextual Clue Aggregation (CCA) module also grants relatively high attentive weights to some words in other arguments, including explosive, Dozens, and Kabul. This implies that these argument words offer crucial information for predicting the role of Islamic State. This visualization illustrates how our Contextual Clue Aggregation (CCA) module effectively identifies relevant contextual clues for candidate spans and models the interaction of information among interconnected arguments within an event.
Figure 8: We provide visualizations of attention weights directed towards the context for various candidate spans within an event.
Additionally, our investigation extends to examining the attention weights associated with various span-trigger pairings within a single event. In Figure 8, we display a heat map created from the attention weights of 30 randomly chosen candidate spans in an event, with respect to their contextual relevance. The heat map reveals that different candidate arguments pay attention to different contextual information, suggesting that our Contextual Clues Aggregation (CCA) module is capable of dynamically identifying and utilizing pertinent contextual clues specific to each candidate argument span. This adaptability highlights the module’s efficiency in tailoring its focus to suit the unique contextual requirements of each argument span within an event.
Figure 9: The visualization of cosine similarity between role representations of two examples from RAMS in sub-figure (a) and MLEE dataset in sub-figure(b) respectively.
Figure 10: Here’s an instance of a t-SNE visualization sourced from the RAMS dataset. The representations of roles and arguments originate from five distinct documents. In sub-figure (a), we utilize average pooling representations encoded by BERT, while in (b), we visualize the representations fused with latent role embeddings.
5.5 Analysis of Role Correlations
To verify the capability of our model in capturing semantic correlations between roles, we visualize the cosine similarity among latent role representations from events within the RAMS and MLEE datasets. Figure 9 presents the results, where roles such as origin and destination, attacker and target, ToLoc and FromLoc exhibit similar representations, indicating their semantic similarity. This visualization validates our model’s efficacy in capturing semantic correlations among roles.
Additionally, to substantiate the effective guidance provided by role representations, we utilize t-SNE visualization JMLR:v9:vandermaaten08a to display arguments corresponding to two distinct roles that appear together in five diverse documents, alongside their respective latent role embeddings. Figure 10a illustrates that arguments associated with the same role but in distinct documents are dispersed throughout the embedding space. This dispersion is attributed to variations in the target events and their contextual settings. However, after integrating latent role embeddings, the representations of arguments linked to identical roles, such as victim or place, cluster more closely. This shows that the Role-based Latent Information Guidance (RLIG) module effectively provides crucial role-specific information. The visualization demonstrates that the RLIG module enhances the alignment of arguments within the same role across different contexts, thereby affirming its utility in providing insightful guidance of roles.
5.6 Complexity and Efficiency
Our CARLG method combines computational efficiency with low complexity while both CCA and RLIG introduce few parameters. As shown in Table 6, for ${\text{CARLG}}_{{\text{large}}}$, CCA utilizes the pre-trained encoder’s attention heads and performs multiplication and normalization operations, adding only about 0.28% new parameters. RLIG introduces approximately 0.81% new parameters totally. This makes our parameter quantity approximately equivalent to the transformer-based encoder plus a MLP classification layer.
Table 7 reports the efficiency comparison of various span-based models on MLEE dataset. Our ${\text{CARLG}}_{{\text{base}}}$ is approximately 2 times faster than ${\text{TSAR}}_{{\text{base}}}$ and can save up to 30.9% of inference time with smaller model size. Additionally, compared to Two-Step, our CARLG only increases 15.9% inference time and 8% model parameters. Such high efficiency and performance makes CARLG a promising choice for practical applications of document-level EAE.
Table 7: Time cost (in seconds) of different span-based models on MLEE dataset. Experiments are run for one epoch on one Tesla A100 GPU.
Model Params Training Time Inference Time
Two-Step 111.96M 229.68 55.29
${\text{TSAR}}_{{\text{base}}}$ 142.03M 465.84 92.79
${\text{CARLG}}_{{\text{base}}}$ 120.87M 280.08 64.08
5.7 Error Analysis
Table 8: Error analysis on the RAMS test set. The best results are in bold font.
Model Wrong Span Over-extract Partial Overlap Wrong Role Total
Two-Step 86 64 47 32 46 275
${\text{TSAR}}_{{\text{base}}}$ 81 48 57 28 19 233
${\text{CARLG}}_{{\text{base}}}$ 79 45 55 29 15 223
Figure 11: An extraction case from RAMS test set base on ${\text{Roberta}}_{{\text{large}}}$. This case shows both CCA module and RLIG module contribute to the argument extraction.
To conduct a more comprehensive examination of the errors made by various models and analyze the underlying reasons for these errors , we manually compared the predictions of 200 randomly selected examples from the RAMS test set against their golden annotations. During our analysis, we categorized the errors into five main categories following xu2022two . Wrong Span refers to the case where a specific role is assigned to a span that does not overlap with the corresponding golden span. We observe that this type of errors often occur when negative words such as ”not” are involved or when there are coreference spans related to the golden span. The category labeled Over-extract pertains to instances where the model predicts an argument role that does not genuinely appear in the document. Some of the extracted spans are either sub-strings of the golden spans (Partial) or exhibit overlaps with them (Overlap). The ”Partial” error category is also commonly observed when there is punctuation, such as a comma, within the golden span. Wrong Role refers to the model correctly identifies the golden span, but assigns the wrong argument role to it. We compare the errors of some span-based methods, that is, Two-Step and ${\text{TSAR}}_{{\text{base}}}$ and ${\text{CARLG}}_{{\text{base}}}$. As shown in Table 8, compared with Two-Step and ${\text{TSAR}}_{{\text{base}}}$, our CARLG model decreases the number of errors from 275 to 223 and effectively reduces Wrong Span errors, which can be attributed to the integration of contextual clues and latent role guidance.
Method RAMS
Arg-I Arg-C
text-davinci-003 46.1 39.5
gpt-3.5-turbo 38.3 31.5
gpt-4 50.4 42.8
CARLG (Ours) 56.71 52.82
Table 9: Comparison with large language model methods on the test set of RAMS.
5.8 Comparison with Large Language Models
Large language models (LLMs) have become a focal point of interest among researchers, given their remarkable versatility across a diverse range of Information Extraction (IE) tasks xu2023large ; li2023semi ; zhang2024ultra . In this paper, we undertake a comprehensive comparison with some large language models on the document-level EAE task. Due to the high cost associated with using GPT-4, its evaluation is limited solely to the RAMS dataset. The experimental results are detailed in Table 9. We utilize the prompts defined in the work zhou2023heuristics and experiments are conducted utilizing three prominent large language models: text-davinci-003 ouyang2022training , gpt-3.5-turbo ouyang2022training and GPT-4 achiam2023gpt , all of which are accessed via OpenAI’s public APIs  [^4] https://openai.com/api/. As illustrated in Table 9, LLMs exhibit a notable performance disparity compared to supervised learning models in the EAE task. Furthermore, the operational costs associated with large models are inherently substantial. In contrast, our approach demonstrates superior efficiency, cost-effectiveness, and overall effectiveness in the document-level EAE task when compared to LLMs.
5.9 Case Study
In order to explore how the CCA and RLIG modules work in actual cases, we conduct the case study for our CARLG. As shown in Figure 11, equipped with CCA and RLIG modules, our CARLG can correctly predict all arguments, while TSAR fails to predict arguments attack. When removing the CCA module, our CARLG fails to predict the argument Talha Hesham Abdo, indicting the CCA module contributes to the argument extraction by providing essential contextual clues. Additionally, when the RLIG module is removed, it fails to predict the argument for the crim role and mispredicts the argument for the jailer role. This can be attributed to that the RLIG module effectively captures the semantic correlations among the roles jailer, detainee, and crim and provides valuable information guidance, allowing the model to make more accurate predictions.
6 Conclusion
In this paper, we introduce the CARLG model for document-level event argument extraction (EAE). CARLG comprises two novel modules: Contextual Clues Aggregation (CCA) and Role-Based Latent Information Guidance (RLIG). CCA effectively aggregates non-argument clue words to capture beneficial contextual information, while RLIG provides guidance based on latent role representations that capture semantic correlations between some roles. Diverse experiments show that our CARLG model surpasses current SOTA EAE models on three large-scale document-level datasets.
While our CARLG model has exhibited remarkable performance in document-level event argument extraction (EAE) tasks with known event triggers, we acknowledge its limitation in depending on the availability of event triggers. In practical situations, event triggers may not always be supplied or known in advance. Therefore, in our future research, we intent to transfer our approach to encompass other EAE tasks that do not rely on explicit trigger words. We aim to conduct extensive experiments to assess its performance in such scenarios. Furthermore, we intend to explore the applicability of the CARLG framework in other information extraction tasks, such as relation extraction.

### v4

@@
-Utilizing Contextual Clues and Role Correlations for Enhancing Document-level Event Argument Extraction
+mode = title
 Abstract
-Document-level event argument extraction (EAE) is a vital but challenging task in the domain of information extraction.
-Most existing approaches focus on the interaction between arguments and event triggers, ignoring two critical points: the information of contextual clues and the semantic correlations among argument roles.
-In this paper, we propose the CARLG (Contextual Clues Aggregation and Role-based Latent Information Guidance) model, which consists of two modules: Contextual Clues Aggregation (CCA) and Role-based Latent Information Guidance (RLIG), effectively leveraging contextual clues and role correlations for improving document-level EAE.
-The CCA module adaptively captures and integrates contextual clues by utilizing context attention weights from a pre-trained encoder.
-The RLIG module captures semantic correlations through role-interactive encoding and provides valuable information guidance with latent role representation.
-Notably, our CCA and RLIG modules are compact, transplantable and efficient, which introduce no more than 1% new parameters and can be easily equipped on other span-base methods with significant performance boost.
-Extensive experiments on the RAMS, WikiEvents, and MLEE datasets demonstrate the superiority of the proposed CARLG model.
-It outperforms previous state-of-the-art approaches by 1.26 F1, 1.22 F1, and 1.98 F1, respectively, while reducing the inference time by 31%.
-Furthermore, we provide detailed experimental analyses based on the performance gains and illustrate the interpretability of our model.
+Document-level event argument extraction is a crucial yet challenging task within the field of information extraction.
+Current mainstream approaches primarily focus on the information interaction between event triggers and their arguments, facing two limitations: insufficient context interaction and the ignorance of event correlations.
+Here, we introduce a novel framework named CARLG (Contextual Aggregation of Clues and Role-based Latent Guidance), comprising two innovative components: the Contextual Clues Aggregation (CCA) and the Role-based Latent Information Guidance (RLIG).
+The CCA module leverages the attention weights derived from a pre-trained encoder to adaptively assimilate broader contextual information, while the RLIG module aims to capture the semantic correlations among event roles.
+We then instantiate the CARLG framework into two variants based on two types of mainstream EAE approaches.
+Notably, our CARLG framework introduces less than 1% new parameters yet significantly improving the performance.
+Comprehensive experiments across the RAMS, WikiEvents, and MLEE datasets confirm the superiority of CARLG, showing significant superiority in terms of both performance and inference speed compared to major benchmarks.
+Further analyses demonstrate the effectiveness of the proposed modules.
 keywords:
-Event Argument Extraction (EAE), Information Extraction (IE), Natural Language Processing (NLP), Natural Language Understanding (NLU).
-journal: Knowledge-Based Systems
+Event Extraction, Information Extraction, Feature Engineering, Natural Language Processing, Knowledge Graph.
 1 Introduction
-Within the domain of Information Extraction (IE), document-level event argument extraction (EAE) is a vital task focused on identifying event arguments and their corresponding roles within textual content associated with specific events.
-This task is a key process of event extraction (EE) xu2023multi ; lu2022explainable ; lv2022trigger and has broad implications across various fields.
-For instance, it is instrumental in the analysis of social media networks mitchell1974social ; DING2023110687 ; YANG2022109939 , enhancing recommendation systems  lietal2020gaia ; LI2024111124 ; DONG2022108954 , and improving dialogue systems  Zhang2020.08.03.20167569 ; ZHAO2023110927 .
+Within the domain of Information Extraction, document-level event argument extraction (EAE) is a vital task focused on identifying event arguments and their corresponding roles within textual content associated with specific events.
+This task is a key process of event extraction (Xu et al., 2023b; Lu et al., 2022; Lv et al., 2022) and has broad applications across various fields.
+For instance, it is instrumental in the analysis of social media networks (Mitchell, 1974; Ding et al., 2023; Yang et al., 2022), enhancing recommendation systems  (Li et al., 2020, 2024a; Dong et al., 2022), and improving dialogue systems  (Zhang et al., 2020a; Zhao et al., 2023).
 As illustrated in Fig.
 1, with the event trigger “homicide” of event type Violence, the goal of EAE system is to identify argument “Freddie Gray” as the role victim, “Baltimore” as the place, and “Officer Caesar Goodson Jr.” as the killer.
-Compared with sentence-level EAE liuetal2018jointly ; waddenetal2019entity ; tongetal2020improving , document-level EAE faces more complex issues, notably long-range dependencies and the need for inferences across multiple sentences, discussed in research ebner2020multi and lietal2021document .
-To solve these problems, some prompt-based methods ma2022prompt ; hsu2023ampere design suitable templates for long text, which solves the long-distance dependency problem and extract arguments in a generative slot-filling manner.
-Additionally, some works utilize graph structures employing heuristic approaches zhang2020two ; pouranbenveysehetal-2022-document , syntactical frameworks xuetal2021document ; liu2022document ; 9204800 , and Graph Neural Networks kipfsemi for the facilitation of reasoning across sentences.
+Compared with sentence-level EAE (Liu et al., 2018; Wadden et al., 2019; Tong et al., 2020), document-level EAE faces more complex issues, notably long-range dependencies and the need for inferences across multiple sentences, as discussed in research (Ebner et al., 2020) and (Li et al., 2021).
+To solve these problems, some prompt-based methods (Ma et al., 2022a; Hsu et al., 2023) design suitable templates for long text, which solves the long-distance dependency problem and extract arguments in a generative slot-filling manner.
+Additionally, some works utilize graph structures employing heuristic approaches (Zhang et al., 2020d; Pouran Ben Veyseh et al., 2022), syntactical frameworks (Xu et al., 2021a; Liu et al., 2022a; Wu et al., 2022), and Graph Neural Networks (Kipf & Welling, ) for the facilitation of reasoning across sentences.
 However, current mainstream approaches ignore two essential aspects: (a) the contextual clue information; (b) the semantic correlations among roles.
+Figure 1: A document from RAMS (Ebner et al., 2020) dataset.
+The event violence has the trigger of Homicide and there are three distinct arguments, each assuming different roles.
+We use the color red to emphasize specific words, representing contextual clues that are significant for the extraction of arguments.
 Contextual clues refer to text elements in the document, excluding the target arguments themselves, which provide beneficial guidance for predicting complex argument roles.
@@
 By considering these contextual clues, the model can leverage wider context information to make more informed and accurate predictions.
-However, many prior studies xu2022two ; lietal2021document ; liu-etal-2023-enhancing-document just rely on the pre-trained transformer encoder to capture global context information in text implicitly, ignoring the necessity of context clues information that is highly relevant to the entity zhou2021document and target event ebner2020multi .
+However, many prior studies (Xu et al., 2022b; Li et al., 2021; Liu et al., 2023b) just rely on the pre-trained transformer encoder to capture global context information in text implicitly, ignoring the necessity of context clues information that is highly relevant to the entity (Zhou et al., 2021) and target event (Ebner et al., 2020).
 Therefore, in this paper, we propose a Contextual Clues Aggregation (CCA) module, which leverages the contextual attention product from the pre-trained transformer encoder to aggregate the event-specific contextual clues for each argument, enhancing the candidate argument representation by incorporating broader and relevant contextual information.
-Figure 1: A document from RAMS ebner2020multi dataset.
-The event violence has the trigger of Homicide and there are three distinct arguments, each assuming different roles.
-We use the color red to emphasize specific words, representing contextual clues that are significant for the extraction of arguments.
-Within the context of real-world document-level EAE task, numerous roles exhibit semantic correlations.
+Figure 2: The visualization represents the co-occurrence frequencies of the top 15 common roles within the RAMS test set, which is defined as the ratio between the co-occurrence times and the sum of the total occurrence times of two roles, and the self-co-occurrence count is set to zero.
+For real-world document-level EAE task, numerous event roles exhibit semantic correlations.
 For example, as illustrated in Figure 1, there exists a strong semantic correlation between the roles of killer and victim, serving as beneficial guidance for extracting arguments associated with these roles within the target event Violence.
@@
 Subsequently, the latent role embeddings are combined with candidate argument spans via pooling and concatenation operations.
-Finally, we achieve alignment of argument spans, triggers, and roles within a unified space through a novel alignment mechanism utilizing Tucker Decomposition, tucker1964extension , which provides valuable information guidance for document-level EAE.
-In summary, to better integrate and leverage our two techniques, we improve existing span-based baselines zhang2020two ; DBLP:journals/corr/abs-1904-05255 for document-level EAE, which has small parameter size and high efficiency.
-Then the CARLG (Contextual Clues Aggregation and Role-based Latent Information Guidance) model is proposed, by equipping CCA and RLIG modules on our baseline for the aforementioned two crucial points respectively.
-Significantly, each module leverages the fine-organized attentive weights obtained from the transformer-based pre-trained model, introducing less than 1% of new parameters.
-Additionally, these two modules can be easily applied to other span-based methods with transformer encoders and provide a significant performance boost in document-level EAE tasks (as shown in our Sec.
-5.2), which are compact and transplantable.
-The computational cost of our model is also lower compared with the recent state-of-the-art approach TSAR, saving no less than 31% of inference time.
-ur contributions are summarized as follows:
-  • 1.
-    An enhanced span-based baseline for document-level EAE tasks is introduced, which demonstrates both parameter and computation efficiency, significantly outperforming existing span-based baselines
-  • 2.
-    We propose a Contextual Clues Aggregation (CCA) module, which adaptively identifies and aggregates the contextual clue information and enhances the representation of candidate arguments by incorporating event-specific contextual information.
-  • 3.
-    We introduce a Role-based Latent Information Guidance (RLIG) module designed to offer guidance based on latent role information, capturing the semantic correlations among roles.
-  • 4.
-    Extensive experiments on the public RAMS and WikiEvents lietal2021document datasets demonstrate that the proposed approach combined effectively with various span-based document-level EAE models and lead to significant performance improvements.
-Our CARLG consistently improves EAE performance on the biological dataset MLEE 10.1093/bioinformatics/bts407 , showing the generality of our approach in different domains.
-Figure 2: The visualization represents the co-occurrence frequencies of the top 15 common roles within the RAMS test set, which is defined as the ratio between the co-occurrence times and the sum of the total occurrence times of two roles, and the self-co-occurrence count is set to zero.
+Finally, we achieve alignment of argument spans, triggers, and roles within a unified space through a novel alignment mechanism utilizing Tucker Decomposition (Tucker et al., 1964), which provides valuable information guidance for document-level EAE.
+In summary, in this paper, we propose the CARLG framework (Contextual Clues Aggregation and Role-based Latent Information Guidance), consisting of CCA module and RLIG module.
+Each module leverages the fine-organized attentive weights obtained from the transformer-based pre-trained model, introducing less than 1% of new parameters.
+We instantiate the CARLG framework into two variants: the span-based $\text{CARLG}_{\text{span}}$ and the prompt-based $\text{CARLG}_{\text{prompt}}$, based on two types of current mainstream EAE approaches.
+Both variants demonstrate significant superiority over leading benchmarks, in terms of both performance and inference speed.
+Our contributions are summarized as follows:
+  • To capture the contextual clues within the text, we propose a Contextual Clues Aggregation (CCA) module.
+It adaptively identifies and aggregates the contextual clue information and enhances the representation of candidate arguments, achieving sufficient context interaction for document-level EAE.
+  • We propose a Role-based Latent Information Guidance (RLIG) module to capture event role correlations, which employs learnable latent role embeddings to enhance the semantic interaction among event roles.
+  • We instantiate the CARLG framework into two variants: the $\text{CARLG}_{\text{span}}$ and the $\text{CARLG}_{\text{prompt}}$, based on two types of current mainstream EAE approaches, which shows the adaptability and transferability of our framework.
+  • Extensive experiments on the public RAMS and WikiEvents (Li et al., 2021) and MLEE (Pyysalo et al., 2012) datasets demonstrate that both CARLG variants show significant superiority in both performance and efficiency.
+Comprehensive ablation studies and analyses illustrate the effectiveness of our proposed framework.
 2 Related Works
@@
 2.1 Event Argument Extraction
-Event argument extraction (EAE) is pivotal for improving event semantic understanding by identifying event arguments within a context, which constitutes a fundamental step in various downstream applications, including question answering liu2022document ; 6544211 ; 9257099 , text summarization allahyari2017text ; 8949697 , and assisted decision-making zhang2020effect ; 9339934 .
+Event argument extraction (EAE) is pivotal for improving event semantic understanding by identifying event arguments within a context, which constitutes a fundamental step in various downstream applications, including question answering (Liu et al., 2022a; Sun et al., 2013; Liu et al., 2022c), text summarization (Allahyari et al., 2017; Huang et al., 2021), and assisted decision-making (Zhang et al., 2020b; Sakai & Iiduka, 2022).
 Previous approaches primarily concentrate on identifying the event trigger and its associated arguments from individual sentences.
-In chenetal2015event , a neural pipeline model was first proposed for event extraction, which was later extended by nguyengrishman2015event ; liuetal2017exploiting to incorporate recurrent neural networks (RNNs) and convolutional neural networks (CNNs) into the pipeline model.
-liuetal2018jointly ; yanetal2019event utilize dependency trees to construct semantic and syntactic dependencies and relations, which captures the dependency of words in a sentence.
-waddenetal2019entity explores all possible spans and utilizes graph neural networks to construct span graphs for information propagation.
-Additionally, transformer-based pre-trained models  waddenetal2019entity ; tongetal2020improving ; liuetal2022dynamic achieve impressive performance in event extraction.
-However, in real-world situations, elements of an event are often expressed across multiple sentences, leading to the emergence of document-level event extraction (DEE) 6910247 ; 7347400 ; SU2024111180 ; SCABORO2023110675 .
-DEE seeks to identify event arguments throughout whole documents, addressing the issue of extended dependencieswangetal2022query ; xu2022two .
+In (Chen et al., 2015), a neural pipeline model was first proposed for event extraction, which was later extended by (Nguyen & Grishman, 2015; Liu et al., 2017) to incorporate recurrent neural networks (RNNs) and convolutional neural networks (CNNs) into the pipeline model.
+(Liu et al., 2018; Yan et al., 2019) utilize dependency trees to construct semantic and syntactic dependencies and relations, which captures the dependency of words in a sentence.
+(Wadden et al., 2019) explores all possible spans and utilizes graph neural networks to construct span graphs for information propagation.
+Additionally, transformer-based pre-trained models  (Wadden et al., 2019; Tong et al., 2020; Liu et al., 2022b) achieve impressive performance in event extraction.
+However, in real-world situations, elements of an event are often expressed across multiple sentences, leading to the emergence of document-level event extraction (DEE) (Cai & Chen, 2015; Jiang et al., 2023; Su et al., 2024; Scaboro et al., 2023).
+DEE seeks to identify event arguments throughout whole documents, addressing the issue of extended dependencies(Wang et al., 2022; Xu et al., 2022b).
 We summarize these recent DEE works into 3 categories: (1) Tagging-based methods.
-Works like wangetal2021cleve ; ducardie2020document apply the BiLSTM-CRF framework for document-level event extraction (DEE).
-Zheng zhengetal2019doc2edag introduces a mixed graph and a tracking module to capture the interrelations between events.
+Works like (Wang et al., 2021; Du & Cardie, 2020a) apply the BiLSTM-CRF framework for document-level event extraction.
+Zheng (Zheng et al., 2019) introduces a mixed graph and a tracking module to capture the interrelations between events.(Li et al., 2021) approaches the issue through conditional generation, whereas (Du et al., 2021) views it as a sequence-to-sequence endeavor.
+Additionally, Wei (Wei et al., 2021) redefines the task in terms of a reading comprehension challenge.
 (2) Span-based methods.
-yangetal2021document proposes an encoder-decoder framework for parallel extraction of structured events.
-Works ebner2020multi ; zhang2020two predict argument roles for selected text segments under a maximum length restriction.
-Furthermore, Xu xu2022two suggests a dual-stream encoder complemented by an AMR-guided graph to manage extended dependencies.
-(3) generation-based methods.
-lietal2021document approaches the issue through conditional generation, whereas duetal2021grit views it as a sequence-to-sequence endeavor.
-Additionally, Wei weietal2021trigger redefines the task in terms of a reading comprehension challenge.
-Furthermore, Ren renetal2022clio integrates argument roles into document encoding to account for multiple role details for nested arguments.
-ma2022prompt ; liu2022document ; XU2023110375 propose a prompt-based method that utilizes slotted prompts for generative slot-filling to extract arguments.
+(Yang et al., 2021) proposes an encoder-decoder framework for parallel extraction of structured events.
+Works (Ebner et al., 2020; Zhang et al., 2020d) predict argument roles for selected text segments under a maximum length restriction.
+Furthermore, Xu (Xu et al., 2022b) suggests a dual-stream encoder complemented by an AMR-guided graph to manage extended dependencies.
+(3) Prompt-based methods.
+Ren (Ren et al., 2022) integrates argument roles into document encoding to account for multiple role details for nested arguments.
+(Ma et al., 2022a; Liu et al., 2022a; Xu et al., 2023c) propose a prompt-based method that utilizes slotted prompts for generative slot-filling to extract arguments.
+Among them, span-based and prompt-based methods are current mainstream methods, which have been demonstrated superior generalizability and competitive performance (Hsu et al., 2023).
 Most existing document-level EAE methods focus on the interaction between arguments and event triggers, ignoring the correlations of roles among events, or the clue information provided by non-arguments context, which can be crucial for the task.
@@
 2.2 Contextual Clues
-Contextual clues refer to the information and cues present in the surrounding context of a given entity, event, or situation clark2016contextual ; zhang2020contextual ; liu2022document .
-These clues provide additional context and can help in understanding the meaning, significance, or relationships associated with the target entity or event peters2018deep .
-NC-DRE zhang2022nc introduces an encoder-decoder structure, leveraging a pre-trained encoding model and a graph neural network (GNN) for decoding.
+Contextual clues refer to the information and cues present in the surrounding context of a given entity, event, or situation (Jiang et al., 2022; Zhang et al., 2020c; Liu et al., 2022a; Li et al., 2024b).
+These clues provide additional context and can help in understanding the meaning, significance, or relationships associated with the target entity or event (Peters et al., 2018).
+NC-DRE (Zhang & Cheng, 2022) introduces an encoder-decoder structure, leveraging a pre-trained encoding model and a graph neural network (GNN) for decoding.
 This setup facilitates the use of non-entity clues in document-level relation extraction.
-Concurrently, a method for localized context pooling zhou2021document has been put forward, allowing for the direct application of attention mechanisms from pre-trained language models.
+Concurrently, a method for localized context pooling (Zhou et al., 2021) has been put forward, allowing for the direct application of attention mechanisms from pre-trained language models.
 This method efficiently leverages the contextual information tied to particular entity pairs, aiding in capturing the beneficial context clues for determining relations.
-Xu xu2021discriminative uses the self-attention mechanism vaswani2017attention to learn a document-level context representation (DLCRep) which contributes to the improvement of information extraction.
+Xu (Xu et al., 2021b) uses the self-attention mechanism (Vaswani et al., 2017) to learn a document-level context representation which contributes to the improvement of information extraction.
 Therefore, contextual clues contain important semantic information, which can provide information guidance for deep mining of semantic associations between entities.
-In this paper, we are the first to incorporate contextual clues techniques into the event argument extraction task and propose the Contextual Clues Aggregation (CCA) module to explicitly capture contextual clues.
+In this paper, we are the first to incorporate contextual clues techniques into the event argument extraction task and propose the Contextual Clues Aggregation module to explicitly capture contextual clues.
 2.3 Relation Embedding
-Many methods in the field of information extraction leverage relation embeddings to enhance the effectiveness of various tasks 9086164 ; 9694521 ; 9123572 .
-CRE chen-badlani-2020-relation presents a novel approach to relation extraction by utilizing sentence representations as relation embeddings, which are assessed through entity embeddings trained concurrently with the knowledge base.
-xuetal2022emrel constructs relation representation explicitly and interacts it with entities and context, which achieves excellent results in multi-triple relation extraction.
-han2022documentlevel employs relation embeddings as a medium and introduces two sub-tasks focused on co-occurrence prediction, aiming to enhance the retrieval of relational facts.
+Many methods in the field of information extraction leverage relation embeddings to enhance the effectiveness of various tasks (Ji et al., 2022; Sun et al., 2023; Lu & Chow, 2022).
+CRE (Chen & Badlani, 2020) presents a novel approach to relation extraction by utilizing sentence representations as relation embeddings, which are assessed through entity embeddings trained concurrently with the knowledge base.
+(Xu et al., 2022a) constructs relation representation explicitly and interacts it with entities and context, which achieves excellent results in multi-triple relation extraction.
+(Han et al., 2022) employs relation embeddings as a medium and introduces two sub-tasks focused on co-occurrence prediction, aiming to enhance the retrieval of relational facts.
 Relation embedding is typically utilized in relation extraction tasks.
@@
 Therefore, we aim to use contextual clues and role correlations embedded in latent embedding to strengthen the document-level EAE performance.
-3 Enhanced Span-based Baseline
-To better integrate our two techniques, we first propose an enhanced span-based baseline for document-level EAE, which exhibits small parameter size and high efficiency.
-We clarify the preliminaries of the document-level EAE task as follows for better understanding our method.
-3.1 Preliminaries
-We follow the definition of event argument extraction in RAMSv1.0 benchmark ebner2020multi . To better understand this task, we first clarify some important notions.
-  • 1.
+Figure 3: Contextual Clues Aggregation Module. The core process of the Role-based Latent Information Guidance algorithm is also similar to this.
+3 Preliminaries
+3.1 Important Notions
+We follow the definition of event argument extraction in RAMSv1.0 benchmark (Ebner et al., 2020). To better understand this task, we first clarify some important notions.
   • Event mention: a text span within a document that mentions an entity object, e.g., the phrase Paul Duggan in Figure 1.
-  • 2.
-    Event trigger (trigger): the key term that most explicitly represents the manifestation of the event, e.g., the word homocide in Figure 1.
-  • 3.
+  • Event trigger (trigger): the key term that most explicitly represents the manifestation of the event, e.g., the word homicide in Figure 1.
   • Event type: the semantic class of an event, e.g., the event type Violence in Figure 1.
-  • 4.
   • Argument role (role): the relation between the event and the argument in which it is involved, for example, the role of Baltimore is identified as Place in Figure 1.
 For this task, argument roles are predetermined for every event type.
-  • 5.
   • Event argument (argument): an entity mention, value or temporal expression playing a specific argument role, e.g., the argument Freddie Gray in Figure 1.
-Figure 3: The architecture or our enhanced baseline.
+3.2 Task Formulation
 Formally, for a document $\mathcal{D}$ composed of $N$ words, denoted as $\mathcal{D}=\{w_{1},w_{2},\ldots,w_{N}\}$, a set of pre-defined event types $\mathcal{E}$, a corresponding set of argument roles $\mathcal{R}_{e}$ associated with each event $e\in\mathcal{E}$, and the event trigger $t$ in $\mathcal{D}$.
 The goal is to detect every pair $(r,s)$ for every event $e$ within the document $\mathcal{D}$, where $r\in\mathcal{R}_{e}$ signifies a role associated with the event $e$, and $s$ denotes a contiguous text span within $\mathcal{D}$.
-In alignment with earlier research xu2022two ; ebner2020multi , we extract event arguments independently for each event within a document.
-3.2 Enhanced Span-based Baseline
-In this section, we improve existing span-based baselines ebner2020multi ; zhang2020two and introduce a new baseline for document-level EAE.
-We integrate other techniques to further improve the performance and the architecture is shown in Figure 3.
-3.2.1 Encoder
-Following xu2022two ; ebner2020multi , we divide a document into several instances, where each instance only contains one event.
-For each document instance $\mathcal{D}$, the target event $e$ and the given trigger word $t_{e}$, we wrap the event trigger with special tokens "*""""*"" * " and regard the embedding of the start "*""""*"" * " token as the trigger embedding, which is adapted from the entity marker technique shi2019simple ; soares2019matching .
-We also add the event type as additional information guidance and the final input sequence $S$ is obtained as follows:
-$S\ =\ \mathrm{[CLS]}\ [E_{e}]\ {e}\ [E_{e}]\ \mathrm{[SEP]}\ w_{1}\ w_{2}*\ t_{e}\ *\ ...\ w_{N}\ \mathrm{[SEP]},$
-where $[E_{e}]$ is the special event token [^1] In our implementation, using [unused] tokens in BERT devlinetal2019bert and introducing specific tokens for RoBERTa DBLP:journals/corr/abs-1907-11692 .
-of event $e$.
-Next, we use a pre-trained transformer encoder to get the embedding of each token as follows:
-$\mathbf{H}=\mathrm{Encoder}(\textsc{$S$}).$ (1)
-Next, we can get the event representation $\mathbf{H}^{e}\in\mathbb{R}^{1\times d}$ for the start $[E_{e}]$ and the context representation $\mathbf{H}^{w}\in\mathbb{R}^{l_{w}\times d}$, where $l_{w}$ represents the length of the word pieces list.
-Notably, when the input sequence is longer than the maximum token length, we employ a dynamic window approach zhou2021document ; zhang2021document to process the long sequence.
-We divide the sequence into overlapping windows and calculate the embeddings for each window.
-Next, we obtain the final representation by averaging the embeddings of tokens that overlap across various windows.
-3.2.2 Candidate Spans Representation
-Filtering Impossible Spans   To eliminate noise information from useless spans, we implement a filtering step to reduce the number of candidate spans.
-Specifically, we exclude argument-impossible spans, such as spans containing commas in the middle.
-For a document $\mathcal{D}$, candidate spans after argument-impossible exclusion can be derived as $\mathcal{S}=\{s_{i:j}\}$.
-This enhancement allows us to eliminate approximately 25% of candidate spans on average, enabling our model to focus on candidate spans that contain valuable information.
-Candidate Spans Representation  For a candidate argument span ranging from $w_{i}$ to $w_{j}$, we first obtain its average pooling representation $\textbf{s}_{i:j}$ as follows:
-$\textbf{s}_{i:j}=\frac{1}{j-i+1}\sum_{k=i}^{j}\mathbf{h}_{k}^{w},$ (2)
-where $\mathbf{h}_{k}^{w}$ is the $k^{\mathrm{th}}$ token embedding from $\mathbf{H}^{w}$.
-As we extract arguments at the span level, where the boundaries can be ambiguous, we employ fully connected neural networks to construct start and end representations as follows:
-$\mathbf{H}^{start}=\mathbf{W}^{start}\mathbf{H},$ (3)
-$\mathbf{H}^{end}=\mathbf{W}^{end}\mathbf{H},$ (4)
-where $\mathbf{H}$ represents the hidden representation of the input sequence $S$, and $\mathbf{W}^{start},\mathbf{W}^{end}\in\mathbb{R}^{d\times d}$ are trainable parameters.
-Subsequently, the candidate span representation $\widetilde{\mathbf{s}}_{i:j}$ can be obtained as follows:
-$\widetilde{\mathbf{s}}_{i:j}=\mathbf{W}^{s}[\mathbf{h}^{start}_{i};\textbf{s}_{i:j};\mathbf{h}^{end}_{j}],$ (5)
-where $\mathbf{W}^{s}\in\mathbb{R}^{3d\times d}$ is the trainable parameter and $\mathbf{h}^{start}_{i}$ and $\mathbf{h}^{end}_{j}$ are the $i^{\mathrm{th}}$ and $j^{\mathrm{th}}$ vector of $\mathbf{H}^{start}$ and $\mathbf{H}^{end}$.
-we establish the boundary loss for identifying the start and end positions in accordance with xu2022two :
-$\begin{array}[]{c}\mathcal{L}_{b}=-\sum_{i=1}^{|\mathcal{D}|}[y_{i}^{s}\log P_{i}^{s}+(1-y_{i}^{s})\log(1-P_{i}^{s})\\ +y_{i}^{e}\log P_{i}^{e}+(1-y_{i}^{e})\log(1-P_{i}^{e})],\end{array}$ (6)
-where $y_{i}^{s}$ and $y_{i}^{e}$ denote golden labels and $P_{i}^{s}=\mathrm{sigmoid}(\mathbf{W}_{1}\mathbf{h}^{start}_{i})$ and $P_{i}^{e}=\mathrm{sigmoid}(\mathbf{W}_{2}\mathbf{h}^{end}_{i})$ represent the probabilities of the word $w_{i}$ being predicted as the start or end word of a golden span.
-3.2.3 Classification Module
-Within each event $e$, for each candidate span $s_{i:j}$, we apply a comparison technique mou-etal-2016-natural ; zeng2020double to enhance its representation.
-In detail, we concatenate the span representation $\widetilde{\mathbf{s}}_{i:j}$ with the trigger representation $\mathbf{h}_{t}$, their absolute disparity $|\mathbf{h}_{t}-\widetilde{\mathbf{s}}_{i:j}|$, element-wise product $\mathbf{h}_{t}\odot\widetilde{\mathbf{s}}_{i:j}$, the embedding of the event type $\mathbf{H}^{e}$ and the embedding of the span length $\mathbf{E}_{len}$.
-Finally, through a feed-forward network, we calculate the predicted probability $P(r_{s_{i:j}})$ of the candidate span $s_{i:j}$:
-$\mathbf{I}_{i:j}=[\widetilde{\mathbf{s}}_{i:j};\mathbf{h}_{t};|\mathbf{h}_{t}-\widetilde{\mathbf{s}}_{i:j}|;\mathbf{h}_{t}\odot\widetilde{\mathbf{s}}_{i:j};\mathbf{H}^{e};\mathbf{E}_{len}],$ (7)
-$P(r_{s_{i:j}})=\mathrm{FFN}(\mathbf{I}_{i:j}).$ (8)
-Considering the issue of imbalanced role distribution and the prevalence of negative samples among candidate arguments, we utilize focal loss 8237586 during the training process.
-Focal loss uses hyperparameters $\alpha$ and $\gamma$ to control the weighting and scaling of the loss function, which allows us to emphasize the importance of positive samples and makes the model focus more on useful information.
-$\begin{array}[]{c}\mathcal{L}_{c}=-\sum_{s_{i:j}\in\mathcal{S}}\alpha{{[1-P(r_{s_{i:j}}=y_{s_{i:j}}}))]}^{\gamma}\\ \cdot\log P(r_{s_{i:j}}=y_{s_{i:j}}).\end{array}$ (9)
-3.2.4 Training and Inference
-During the training process, our model extracts single event at a time.
-Our training loss consists of two parts: classification loss $\mathcal{L}_{c}$ and boundary loss $\mathcal{L}_{b}$, and $\lambda$ is hyperparameter.
-$\mathcal{L}=\mathcal{L}_{c}+\lambda\mathcal{L}_{b}.$ (10)
-Consistent with the training process, we predict and choose the argmax role for each candidate span at inference time.
-If more than one argument span is predicted with the same role, we reserve all spans for this role at inference time.
-Figure 4: The main structure of CARLG involves an input sequence, encompassing roles and event types, passing through a transformer-based encoder.
+In alignment with earlier research (Xu et al., 2022b; Ebner et al., 2020), we extract event arguments independently for each event within a document.
+4 CARLG
+In this section, we introduce CARLG framework, comprising Contextual Clues Aggregation (CCA) module which adaptively aggregates contextual clues and Role-based Latent Information Guidance (RLIG) module which effectively captures semantic correlations among roles.
+The core purpose of these two modules is to enhance the representation of the arguments to be predicted.
+Additionally, they can be applied for any approach that relies on argument representation, including span-based methods that make predictions using candidate span representation, and prompt-based methods that predict based on argument slot representation.
+4.1 Contextual Clues Aggregation
+Given that different events focus on different contextual information, the CCA module adaptively aggregates contextual information based on specific event feature to obtain a broader context representation.
+Since the event trigger is crucial for the arguments within an event, we adaptively aggregate context information based on each trigger-argument pair, denoted as $(t,a_{i})$ and $a_{i}$ is the $i^{th}$ argument.
+We leverage the attentive heads from the pre-trained transformer encoder, which enables the adoption of attentions from the pre-trained language model instead of training new attention layers.
+Figure 3 illustrates the process of the proposed CCA module.
+Specifically, we encode the input text sequence $S$ and get the encoded representation $\textbf{H}^{C}$ through a transformer-based pre-trained language model.
+We can obtain the attention head matrix $\textbf{A}^{C}\in\mathbb{R}^{H\times l\times l}$ from the final transformer layer, where $H$ is the number of attention heads and $l$ represents the text sequence length.
+We then extract the contextual attention heads of the trigger and each argument through the pooling operation, denoted as $\textbf{A}^{C}_{t}\in\mathbb{R}^{l}$ and $\textbf{A}^{C}_{a_{i}}\in\mathbb{R}^{l}$.
+It’s important to note that since we do not know in advance what the argument is, the way it is obtained varies by method.
+For span-based methods, we consider the tokens within the span as the argument.
+For prompt-based methods, we treat the slot in the prompt template as the argument.
+We explain this in detail in Section 5.
+Then for each argument, we get the contextual clue information $\textbf{c}_{a_{i}}\in\mathbb{R}^{d}$ by applying multiplication to the attentions and then normalizing the result:
+$\centering\begin{array}[]{c}\textbf{p}^{C}_{a_{i}}=softmax(\textbf{A}^{C}_{a_{i}}\cdot\textbf{A}_{t}^{C}\;)\vspace{1.3ex},\\ \textbf{c}_{a_{i}}={\mathbf{H}^{C}}^{T}\ \textbf{p}^{C}_{a_{i}},\end{array}\@add@centering$ (1)
+where $\cdot$ means dot product and $\textbf{p}^{C}_{a_{i}}\in\mathbb{R}^{l_{w}}$ is the computed contextual attention weight vector.
+$T$ is the transpose symbol.
+$\textbf{c}_{a_{i}}$ is the enhanced context vector specific to the argument $a_{i}$ and trigger $t$.
+The detailed algorithm is shown in Algorithm 1.
+Algorithm 1 Contextual Clues Aggregation
+1:Input text sequence $S$, pre-trained transformer encoder $Encoder$
+2:Context-enhanced argument representations $\textbf{c}_{a_{i}}$
+3:$\textbf{H}^{C}\leftarrow Encoder(S)$triangleright Encode $S$ using the pre-trained transformer
+4:Obtain attention head matrix $\textbf{A}^{C}\in\mathbb{R}^{H\times l\times l}$ from the final layer of $Encoder$
+5:for each argument $a_{i}$ do
+6:     Identify the index $t_{idx}$ of the trigger in the sequence
+7:     Identify the index range $a_{i_{idx}}$ of the argument $a_{i}$ in the sequence
+8:     Extract $\textbf{A}^{C}_{t}\in\mathbb{R}^{l}$ for the trigger by selecting the $t_{idx}$-th row from $\textbf{A}^{C}$
+9:     Extract $\textbf{A}^{C}_{a_{i}}\in\mathbb{R}^{l}$ for the argument $a_{i}$ by aggregating rows in $a_{i_{idx}}$ from $\textbf{A}^{C}$
+10:     Compute contextual attention weights: $\textbf{p}^{C}_{a_{i}}=\text{softmax}(\textbf{A}^{C}_{a_{i}}\cdot\textbf{A}_{t}^{C})$
+11:     Obtain context vector: $\textbf{c}_{a_{i}}={\mathbf{H}^{C}}^{T}\textbf{p}^{C}_{a_{i}}$
+12:end for
+13:return $\textbf{c}_{a_{i}}$ for each argument $a_{i}$
+4.2 Role-based Latent Information Guidance
+In this section, we propose Role-based Latent Information Guidance (RLIG) module, consisting of Role-interactive Encoding and Role Information Fusion.
+4.2.1 Role-interactive Encoding
+In NLP applications, the inclusion of prefixes, like relation and other injected information, significantly enhances the performance of downstream tasks (Xu et al., 2022a; Hsu et al., 2023; Liu et al., 2023a).
+To effectively capture the semantic correlations between roles, we integrate the information of role types as prefixes into the input and utilize multi-head attention to facilitate interaction between the context and roles.
+We generate latent role embeddings by assigning unique special tokens   [^1] In our implementation, using [unused] tokens in BERT (Devlin et al., 2019) and introducing specific tokens for RoBERTa (Liu et al., 2019) to represent different roles in the framework of a pre-trained model, with each role type gaining a unique latent representation.
+Considering that role names carry valuable semantic information (Wang et al., 2022), we encase role names in special tokens indicative of their role types, selecting the embedding of the initial special token as the role’s representation.
+This approach ensures the capture of unique attributes and subtle semantic distinctions of each role.
+For instance, the role “Victim” is denoted as $[R_{0}]\quad Victim\quad[R_{0}]$, with $[R_{0}]$ representing the specialized token for “Victim”.
+Algorithm 2 Role-based Latent Information Guidance
+1:Input text sequence $S$, set of roles $\{r_{0},r_{1},\ldots,r_{n}\}$, pre-trained transformer encoder
+2:Role-enhanced argument representations $\textbf{r}_{a_{i}}$
+3:Enhance $S$ by concatenating role type elements:
+4:$S^{\prime}=[R_{0}]\ {r_{0}}\ [R_{0}]\ [R_{1}]\ {r_{1}}\ [R_{1}]\ ...\ \mathrm{[SEP]}\ w_{1}\ w_{2}\ ...\ w_{N}$
+5:Encode $S^{\prime}$ using the pre-trained transformer to obtain $\mathbf{H}^{C}$ and $\mathbf{H}^{R}$
+6:for each argument $a_{i}$ do
+7:     Identify indices $t_{idx}$ and $a_{i_{idx}}$ for trigger $t$ and argument $a_{i}$ in $S^{\prime}$
+8:     Extract $\textbf{A}^{R}_{t}\in\mathbb{R}^{l_{r}}$ by selecting the attention vector corresponding to
+9:  $t_{idx}$ from $\mathbf{H}^{R}$
+10:     Extract $\textbf{A}^{R}_{a_{i}}\in\mathbb{R}^{l_{r}}$ by selecting the attention vector corresponding to
+11:  $a_{i_{idx}}$ from $\mathbf{H}^{R}$
+12:     Compute role attention weights: $\textbf{p}^{R}_{a_{i}}=\text{softmax}(\textbf{A}^{R}_{a_{i}}\cdot\textbf{A}^{R}_{t})$
+13:     Obtain role-specific information: $\textbf{r}_{a_{i}}={\mathbf{H}^{R}}^{T}\textbf{p}^{R}_{a_{i}}$
+14:end for
+15:return $\textbf{r}_{a_{i}}$ for each argument $a_{i}$
+Given the input sequence $S$, we enhance it by concatenating role type elements in the following manner:
+$\quad S\ =\ \textsc{ $[R_{0}]\ {r_{0}}\ [R_{0}]\ [R_{1}]\ {r_{1}}\ [R_{1}]\ ...\ \mathrm{[SEP]}$}$
+$\ \quad\quad\quad\quad w_{1}\ w_{2}\ *\ t_{e}\ *\ ...\ w_{N},$
+where $[R_{0}]$ and $[R_{1}]$ denote the special role type tokens for $r_{0}$ and $r_{1}$, respectively, and we utilize the final $\mathrm{[SEP]}$ token to signify the “none” category.
+By employing role-interactive encoding, we are able to get the context representation $\mathbf{H}^{C}\in\mathbb{R}^{l_{w}\times d}$, and the role representation $\mathbf{H}^{R}\in\mathbb{R}^{l_{r}\times d}$.
+Here, $l_{w}$ denotes the length of the list of word pieces, and $l_{r}$ represents the length of the list of roles.
+4.2.2 Role Information Fusion
+Through role-interactive encoding, the role embeddings are able to capture semantic correlations and dynamically adapt to the target event and context.
+To ensure each candidate argument benefits from the guidance of role-specific information, we adapt our method for aggregating contextual clues to selectively incorporate role information.
+The aggregated role information $\textbf{r}_{a_{i}}\in\mathbb{R}^{d}$ for argument ${a_{i}}$ is obtained as follows:
+$\begin{array}[]{c}\textbf{p}^{R}_{a_{i}}=softmax(\textbf{A}^{R}_{a_{i}}\cdot\textbf{A}_{t}^{R}\;)\vspace{1.3ex},\\ \textbf{r}_{a_{i}}={\mathbf{H}^{R}}^{T}\ \textbf{p}^{R}_{a_{i}},\end{array}$ (2)
+where $\textbf{A}^{R}_{a_{i}}\in\mathbb{R}^{l_{r}}$ represents the role attention for argument $a_{i}$, while $\textbf{A}^{R}_{t}\in\mathbb{R}^{l_{r}}$ denotes the role attention associated with the trigger $t$.
+Additionally, the vector $\textbf{p}^{r}_{a_{i}}\in\mathbb{R}^{l_{r}}$ characterizes the computed attention weight vector for roles.
+$\textbf{r}_{a_{i}}$ is the aggregated role vector for argument $a_{i}$.
+The detailed algorithm is shown in Algorithm 2.
+Then the enhanced context vector $\textbf{c}_{a_{i}}$ and role vector $\textbf{r}_{a_{i}}$ can be utilized to enhance the representation e of the argument $a_{i}$ to be predicted.
+5 Adapting CARLG to Span-based and Prompt-based EAE Approaches
+We instantiate the CARLG framework into two variants: the span-based $\text{CARLG}_{\text{span}}$ and the prompt-based $\text{CARLG}_{\text{prompt}}$, based on two types of current mainstream EAE approaches.
+5.1 $\text{CARLG}_{\text{span}}$
+We apply our CARLG framework to span-based EAE approaches, where the architecture is shown in Figure 4.
+Since the application of CARLG framework is consistent across span-based methods, we choose the architecture of TSAR (Xu et al., 2022b) as the foundation to illustrate the span-based baseline [^2] We extract the core architecture of TASR as the span-based baseline, removing the two-stream module and the AMR graph module., and we detail the application of our CARLG framework to other typical span-based methods in the experimental Section 6.2.
+Figure 4: The main structure of $\text{CARLG}_{\text{span}}$.
+The input sequence, encompassing roles and event types, passes through a transformer-based encoder.
 This encoder generates output in the form of context representations, role representations, and attention heads.
@@
 Finally, the classification module leverages this representation to forecast argument roles for all potential spans.
-4 CARLG Model
-In this section, we introduce Contextual Clues Aggregation (CCA) module which adaptively aggregates contextual clues and Role-based Latent Information Guidance (RLIG) module which effectively captures semantic correlations among roles.
-These two tecniques are then well equipped on our proposed baseline and our CARLG model are shown in Figure 4.
-4.1 Contextual Clues Aggregation
-The average pooling representation that previous span-based methods xu2022two ; ebner2020multi ; zhang2020two disregards important contextual information from non-argument words.
-Although the self-attention mechanism of the pre-trained encoder can capture token-level interactions, it lacks specific global interaction tailored to the event and candidate arguments.
-To address this limitation, we propose a Contextual Clues Aggregation (CCA) module that selects and combines relevant contextual information for each tuple comprising a candidate span and the event trigger word, denoted as $(s_{i:j},t)$.
-In this module, as shown in Figure 5, we leverage the attentive heads from the pre-trained transformer encoder, which enables the adoption of attentions from the pre-trained language model instead of training new attention layers.
-We employ the token-level attention heads $\textbf{A}^{w}\in\mathbb{R}^{H\times l_{w}\times l_{w}}$ obtained from the final transformer layer of the pre-trained language model.
+5.1.1 Role-interactive Encoder
+We apply Algorithm 2 and construct a role-interactive encoder to capture role correlations.
+Following (Xu et al., 2022b; Ebner et al., 2020), we divide a document into several instances, where each instance only contains one event.
+For an instance, given the input document $\mathcal{D}$, the target event $e$ and its associated set of roles $\mathcal{R}_{e}=\{{r}_{1},{r}_{2},{r}_{3},...\}$, we enhance the input by concatenating these elements in the following manner:
+$\quad S_{\text{span}}\ =\ \mathrm{[CLS]}\ [E_{e}]\ {e}\ [E_{e}]\ \mathrm{[SEP]}\ w_{1}\ w_{2}\ *\ t_{e}\ *\ ...\ w_{N}$
+$\quad\quad\mathrm{[SEP]}\textsc{ \ $[R_{0}]\ {r_{0}}\ [R_{0}]\ [R_{1}]\ {r_{1}}\ [R_{1}]\ ...\ \mathrm{[SEP]}$},$
+where $[R_{0}]$ and $[R_{1}]$ denote the special role type tokens for $r_{0}$ and $r_{1}$, respectively, and we utilize the final $\mathrm{[SEP]}$ token to signify the “none” category.
+Next, we use a pre-trained transformer encoder to get the embedding of each token as follows:
+$\mathbf{H}=\mathrm{Encoder}(S_{\text{span}}).$ (3)
+We can get the event representation $\mathbf{H}^{e}\in\mathbb{R}^{1\times d}$ for the start $[E_{e}]$ and the context representation $\mathbf{H}^{w}\in\mathbb{R}^{l_{w}\times d}$, where $l_{w}$ represents the length of the word pieces list.
+Notably, when the input sequence is longer than the maximum token length, we employ a dynamic window approach (Zhou et al., 2021; Zhang et al., 2021) to process the long sequence.
+We divide the sequence into overlapping windows and calculate the embeddings for each window.
+Next, we obtain the final representation by averaging the embeddings of tokens that overlap across various windows.
+5.1.2 Contextual Clues Aggregation
+We apply Algorithm 1 to the span-based representations, capturing context clues to enhance the representation of candidate spans.
+We employ the token-level attention heads $\textbf{A}^{C}\in\mathbb{R}^{H\times l_{w}\times l_{w}}$ obtained from the final transformer layer of the pre-trained language model.
 Here, $H$ represents the number of attention heads, and $l_{w}$ signifies the length of word pieces.
 Utilizing these attention heads, we compute the context attention $\textbf{A}^{C}_{i:j}\in\mathbb{R}^{l_{w}}$ for each candidate span spanning from $w_{i}$ to $w_{j}$, by applying average pooling:
-$\textbf{A}^{C}_{i:j}=\frac{1}{H(j-i+1)}\sum_{h=1}^{H}\sum_{m=i}^{j}\mathbf{A}^{w}_{h,m}.$ (11)
-Then for span-trigger tuple $(s_{i:j},t)$, we get the contextual clue information $\textbf{c}^{s_{i:j}}\in\mathbb{R}^{d}$ which is crucial for the candidate span, by applying multiplication to the attentions and then normalizing the result:
-Figure 5: Contextual Clues Aggregation Module.
-$\begin{array}[]{c}\textbf{p}^{c}_{i:j}=softmax(\textbf{A}^{C}_{i:j}\cdot\textbf{A}_{t}^{C}\;)\vspace{1.3ex},\\ \textbf{c}^{s_{i:j}}={\mathbf{H}^{w}}^{T}\ \textbf{p}^{c}_{i:j},\end{array}$ (12)
+$\textbf{A}^{C}_{i:j}=\frac{1}{H(j-i+1)}\sum_{h=1}^{H}\sum_{m=i}^{j}\mathbf{A}^{C}_{h,m}.$ (4)
+Then for span-trigger tuple $(s_{i:j},t)$, we get the contextual clue information $\textbf{c}_{s_{i:j}}\in\mathbb{R}^{d}$ which is crucial for the candidate span, by applying multiplication to the attentions and then normalizing the result:
+$\begin{array}[]{c}\textbf{p}^{C}_{i:j}=softmax(\textbf{A}^{C}_{i:j}\cdot\textbf{A}_{t}^{C}\;)\vspace{1.3ex},\\ \textbf{c}_{s_{i:j}}={\mathbf{H}^{C}}^{T}\ \textbf{p}^{C}_{i:j},\end{array}$ (5)
 where $\textbf{A}^{C}_{t}\in\mathbb{R}^{l_{w}}$ is the contextual attention of trigger $t$, $\cdot$ means dot product and $\textbf{p}^{C}_{i:j}\in\mathbb{R}^{l_{w}}$ is the computed contextual attention weight vector.
 $T$ is the transpose symbol.
-We can apply our CCA module to other span-based methods by modifying Eq. 2 and fusing the contextual clue information $\textbf{c}^{s_{i:j}}$ into the average pooling representation as follows:
-$\textbf{s}^{c}_{i:j}=tanh(\mathbf{W}^{CCA}[\frac{1}{j-i+1}\sum_{k=i}^{j}\mathbf{h}_{k}^{w};\textbf{c}^{s_{i:j}}]),$ (13)
-where $\mathbf{W}^{CCA}\in\mathbb{R}^{2d\times d}$ is learnable parameter. $\textbf{s}^{c}_{i:j}$ is the enhanced representation of the candidate span $s_{i:j}$.
-Figure 6: Role-based Latent Information Guidance Module. Sub-figure (a) illustrates the information fusion and (b) illustrates the Latent Role Information Guidance operation.
-4.2 Role-based Latent Information Guidance
-In this section, we propose Role-based Latent Information Guidance (RLIG) module, which is shown in Figure 6.
-4.2.1 Role-interactive Encoding
-To capture the semantic correlations between roles effectively, we integrate the information of role types into the input and utilize multi-head attention to facilitate interaction between the context and roles.
-This enables us to obtain the context and role representation in a shared knowledge space.
-We generate latent role embeddings by assigning unique special tokens   [^2] In our implementation, using [unused] tokens in BERT devlinetal2019bert and introducing specific tokens for RoBERTa DBLP:journals/corr/abs-1907-11692 to embody different roles in the framework of a pre-trained model, with each role type gaining a unique latent representation.
-Considering that role names carry valuable semantic information wangetal2022query , we encase role names in special tokens indicative of their role types, selecting the embedding of the initial special token as the role’s representation.
-This approach ensures the capture of unique attributes and subtle semantic distinctions of each role.
-For instance, the role “Victim” is denoted as $[R_{0}]\quad Victim\quad[R_{0}]$, with $[R_{0}]$ representing the specialized token for “Victim”.
-Given the input document $\mathcal{D}$, the target event $e$ and its associated set of roles $\mathcal{R}_{e}=\{{r}_{1},{r}_{2},{r}_{3},...\}$, we enhance the input sequence described in Sec.
-3.2.1 by concatenating these elements in the following manner:
-$S\ =\ \mathrm{[CLS]}\ [E_{e}]\ {e}\ [E_{e}]\ \mathrm{[SEP]}\ w_{1}\ w_{2}\ *\ t_{e}\ *\ ...\ w_{N}\ \mathrm{[SEP]}$
-$[R_{0}]\ {r_{0}}\ [R_{0}]\ [R_{1}]\ {r_{1}}\ [R_{1}]\ ...\ \mathrm{[SEP]}$,
-where $[R_{0}]$ and $[R_{1}]$ denote the special role type tokens for $r_{0}$ and $r_{1}$, respectively, and we utilize the final $\mathrm{[SEP]}$ token to signify the ”none” category.
-By employing role-interactive encoding, we are able to get the event representation $\mathbf{H}^{e}\in\mathbb{R}^{1\times d}$ for the start event $[E_{e}]$, the context representation $\mathbf{H}^{w}\in\mathbb{R}^{l_{w}\times d}$, and the role representation $\mathbf{H}^{r}\in\mathbb{R}^{l_{r}\times d}$, all from $\mathbf{H}^{s}$.
-Here, $l_{w}$ denotes the length of the list of word pieces, and $l_{r}$ represents the length of the list of roles.
-4.2.2 Role Information Fusion
-Through role-interactive encoding, the role embeddings are able to capture semantic correlations and dynamically adapt to the target event and context.
+Figure 5: Detailed RLIG module for $\text{CARLG}_{\text{span}}$. Sub-figure (a) illustrates the information fusion and (b) illustrates the Latent Role Information Guidance operation.
+5.1.3 Latent Role Information Guidance
+Since span-based methods ultimately classify candidate spans based on the categories of roles, we propose Role Information Guidance on top of Algorithm 2 to better utilize the learned latent role representations, as shown in Figure 5.
+Latent Role Information Fusion  Through the role-interactive encoding, the role embeddings are able to capture semantic correlations and dynamically adapt to the target event and context.
 To ensure each candidate argument benefits from the guidance of role-specific information, we adapt our method for aggregating contextual clues to selectively incorporate role information.
-The aggregated role information $\textbf{r}^{s_{i:j}}\in\mathbb{R}^{d}$ for ${s_{i:j}}$ is obtained through contextual pooling, which involves adjusting the procedure described in Equations 11 and 12:
-$\begin{array}[]{c}\textbf{A}^{R}_{i:j}=\frac{1}{H(j-i+1)}\sum_{h=1}^{H}\sum_{m=i}^{j}\mathbf{A}^{r}_{h,m}\vspace{1.3ex},\\ \textbf{p}^{r}_{i:j}=softmax(\textbf{A}^{R}_{i:j}\cdot\textbf{A}_{t}^{R}\;)\vspace{1.3ex},\\ \textbf{r}^{s_{i:j}}={\mathbf{H}^{r}}^{T}\ \textbf{p}^{r}_{i:j},\end{array}$ (14)
+The aggregated role information $\textbf{r}_{s_{i:j}}\in\mathbb{R}^{d}$ for ${s_{i:j}}$ is obtained through contextual pooling, which involves adjusting the procedure described in Algorithm 2:
+$\begin{array}[]{c}\textbf{A}^{R}_{i:j}=\frac{1}{H(j-i+1)}\sum_{h=1}^{H}\sum_{m=i}^{j}\mathbf{A}^{R}_{h,m}\vspace{1.3ex},\\ \textbf{p}^{R}_{i:j}=softmax(\textbf{A}^{R}_{i:j}\cdot\textbf{A}_{t}^{R}\;)\vspace{1.3ex},\\ \textbf{r}_{s_{i:j}}={\mathbf{H}^{R}}^{T}\ \textbf{p}^{r}_{i:j},\end{array}$ (6)
 where $\textbf{A}^{R}\in\mathbb{R}^{H\times l_{w}\times l_{r}}$ represents the role attention heads obtained from the final transformer layer of the pre-trained model.
@@
 Additionally, the vector $\textbf{p}^{R}_{i:j}\in\mathbb{R}^{l_{r}}$ characterizes the computed attention weight vector for roles.
-We can modify Eq.
-13 and integrate the latent role information $\textbf{r}^{s_{i:j}}$ into the average pooling representation as follows:
-$\textbf{s}^{r}_{i:j}=tanh(\mathbf{W}^{RLIG}[\frac{1}{j-i+1}\sum_{k=i}^{j}\mathbf{h}_{k}^{w};\textbf{r}^{s_{i:j}}]),$ (15)
-where $\mathbf{W}^{RLIG}\in\mathbb{R}^{2d\times d}$ is learnable parameter.
-4.2.3 Latent Role Information Guidance
+We can fuse the contextual clue information $\textbf{c}_{s_{i:j}}$ and aggregated role information $\textbf{r}_{s_{i:j}}$into the average pooling representation as follows:
+$\textbf{s}_{i:j}=tanh(\mathbf{W}^{Span}[\frac{1}{j-i+1}\sum_{k=i}^{j}\mathbf{h}_{k}^{w};\textbf{c}_{s_{i:j}};\textbf{r}_{s_{i:j}}]),$ (7)
+where $\mathbf{W}^{Span}\in\mathbb{R}^{3d\times d}$ is learnable parameter. $\textbf{s}_{i:j}$ is the enhanced representation of the candidate span $s_{i:j}$.
+Role Information Guidance  To maximize the utilization of the learned role representations, we introduce the Latent Role Information Guidance algorithm, tailored to the span-based architecture.
 We first map the original $d$-dimensional role features into a reduced dimension $d^{\prime}$ via a fully connected network to obtain the latent role information $\mathbf{H}^{r^{\prime}}\in\mathbb{R}^{l_{r}\times d^{{}^{\prime}}}$ as follows:
@@
 where $\mathbf{W}_{3}\in\mathbb{R}^{d\times d^{{}^{\prime}}}$ and $\mathbf{b}_{w}\in\mathbb{R}^{d^{{}^{\prime}}}$ are learnable parameters.
-To fully utilize the expressive power of role embeddings, we propose a factorization-based alignment approach using Tucker decomposition  tucker1964extension .
+To fully utilize the expressive power of role embeddings, we propose a factorization-based alignment approach using Tucker decomposition  (Tucker et al., 1964).
 We define candidate span-trigger-role triples set $\mathcal{T}=\{<s_{i:j},r_{k},t_{e}>|s_{i:j}\in\mathcal{S},r_{k}\in\mathcal{R}_{e}\}$.
-Then we introduce a core tensor $\mathbf{Z}\in\mathbb{R}^{d_{i}\times d^{{}^{\prime}}}$ where $d_{i}$ is the hidden dimension of the combined representation $\mathbf{I}_{i:j}$ calculated in Eq.
-7, and for simplity, the validity for each triple $<s_{i:j},r_{k},t_{e}>\in\mathcal{T}$ is scored as follows:
+Then we introduce a core tensor $\mathbf{Z}\in\mathbb{R}^{d_{i}\times d^{{}^{\prime}}}$ where $d_{i}$ is the hidden dimension of the combined representation $\mathbf{I}_{i:j}$.
+$\mathbf{I}_{i:j}=FFN([\mathbf{h}_{t};\textbf{s}_{i:j}])$, and $FFN$ is the full-connected neural network.
+For simplity, the validity for each triple $<s_{i:j},r_{k},t_{e}>\in\mathcal{T}$ is scored as follows:
 $\varphi(s_{i:j},r_{k},t_{e})\;=\;\sigma(\mathbf{Z}\;\times_{1}\;\mathbf{I}_{i:j}\times_{2}\;{\widehat{\mathbf{r}}}_{\mathrm{k}}\;+b_{k}),$ (9)
 where ${\widehat{\mathbf{r}}}_{\mathrm{k}}=\mathbf{H}^{r^{\prime}}_{k,:}$ represents the transformed role features and ×$\times_{n}$ denotes the tensor product operation along the $n^{th}$ dimension.
-The calculation of $\varphi$ for all triples is performed concurrently via a batched tensor product technique, and the model is trained employing focal loss as described in reference 8237586 .
+The calculation of $\varphi$ for all triples is performed concurrently via a batched tensor product technique, and the model is trained employing focal loss as described in reference (Lin et al., 2017).
+5.1.4 Training and Inference
+Considering the issue of imbalanced role distribution and the prevalence of negative samples among candidate arguments, we utilize focal loss (Lin et al., 2017) during the training process.
+Focal loss uses hyperparameters $\alpha$ and $\gamma$ to control the weighting and scaling of the loss function, which allows us to emphasize the importance of positive samples and makes the model focus more on useful information.
+$\begin{array}[]{c}\mathcal{L}_{c}=-\sum_{s_{i:j}\in\mathcal{S}}\alpha{{[1-P(\varphi(s_{i:j},r_{k},t_{e})\;=y_{s_{i:j}}}))]}^{\gamma}\vspace{1.3ex}\\ \cdot\log P(\varphi(s_{i:j},r_{k},t_{e})\;=y_{s_{i:j}}).\end{array}$ (10)
+Consistent with the training process, we predict and choose the argmax role for each candidate span at inference time.
+If more than one argument span is predicted with the same role, we reserve all spans for this role at inference time.
+5.2 $\text{CARLG}_{\text{prompt}}$
+In this section, we apply our CARLG framework to prompt-based EAE approaches, where the architecture is shown in Figure 7.
+We choose the architecture of PAIE (Ma et al., 2022b) as the foundation to illustrate the prompt-based baseline.
+Here, we only provide a detailed introduction to the parts related to our CCA and RLIG and the other parts are omitted.
+Figure 6: The main structure of $\text{CARLG}_{\text{prompt}}$ involves an input sequence, encompassing roles and prompts, passing through an encoder-decoder architecture.
+This encoder generates output in the form of context representations, role representations, and attention heads.
+The Contextual Clue Aggregation (CCA) module dynamically consolidates contextual clues, while the Role-based Latent Information Generation (RLIG) module generates latent role embeddings and captures role correlations.
+Subsequently, the context vector and the latent role vector are sent to the span selector to predict corresponding spans of each slot.
+Ultimately, based on the predicted spans, we extract the corresponding argument for each role from the text.
+5.2.1 Input Preprocess
+We utilize manual templates to generate the prompt $P$ for each instance [^3] Inspired by this work (Ma et al., 2022b), manual template has the most stable performance and usually the better result than other prompt variants..
+The template for each event type is pre-defined as a sentence with slots $role\right>$.
+The generated prompt for each specific event type is a filled template where slots are substituted with actual role names.
+We use $s_{k}$ to denote the $k$-th slot in the prompt $P$.
+[…]
+The above case shows a concrete example for the prompt creation, and $s_{0}$ = $victim\right>$, $s_{1}$ = $place\right>$, $s_{2}$ = $killer\right>$.
+Next, we concatenate the input text with the prompt. Following Algorithm 2, we prepend the role type to the sequence to obtain the final sequence $S_{\text{prompt}}$.
+5.2.2 Prompt-based CCA and RLIG
+We adopt an encoder-decoder architecture.
+The encoder is employed to encode the input text, while the decoder is tasked with deriving the event-oriented context and context-oriented prompt representation $\mathbf{H}_{\mathrm{de}}$:
+$\begin{array}[]{c}[\mathbf{A};\mathbf{H}_{\mathrm{en}}]=\mathrm{Encoder}(S_{\text{prompt}}),\vspace{1.0ex}\end{array}$ (11)
+where the $\mathrm{Encoder}$ is a pre-trained transformer-based encoder.
+$\mathbf{A}\in\mathbb{R}^{H\times l_{s}\times l_{s}}$ is the multi-head attention matrix and $\mathbf{H}_{\mathrm{en}}$.
+$H$ is the attention head numbers and $l_{s}$ is the length of input sequence $S_{\text{prompt}}$.
+For prompt-based EAE approaches, the representation of event arguments is indicated through slots in the prompt.
+Therefore, we replace the $\textbf{A}^{C}_{a_{i}}$ in Algorithm 1 and $\textbf{A}^{R}_{a_{i}}$ in Algorithm 2 with the corresponding attention heads of the slots in the input sequence, respectively.
+The remaining operations are consistent with the steps in Algorithm 1 and Algorithm 2.
+Ultimately, we can obtain the enhanced context vector $\textbf{c}_{s_{i}}$ and role vector $\textbf{r}_{s_{i}}$ for the the slot of argument $a_{i}$.
+Then we utilize a decoder to derive the event-oriented context and context-oriented prompt representation $\mathbf{H}_{\mathrm{de}}$:
+$\begin{array}[]{c}\mathbf{H}_{\mathrm{de}}=\mathrm{Decoder}(\mathbf{H}_{\mathrm{en}}),\vspace{1.0ex}\end{array}$ (12)
+where the $\mathrm{Decoder}$ is a transformer-based decoder and $\mathbf{H}_{\mathrm{de}}\in\mathbb{R}^{l_{s}\times d}$.
+Next, we integrate $\textbf{c}_{s_{i}}$ and $\textbf{r}_{s_{i}}$ into the decoder output $\mathbf{h}_{s{i}}\in\mathbb{R}^{d}$ for slot $s_{i}$, resulting in an updated representation $\tilde{\mathbf{h}}{s{i}}\in\mathbb{R}^{d}$::
+$\tilde{\mathbf{h}}_{s_{i}}=\mathrm{tanh}(\mathbf{W}^{Prompt}[\mathbf{h}_{s_{i}};\textbf{c}_{s_{i}};\textbf{r}_{s_{i}}]),$ (13)
+where $\mathbf{W}^{Prompt}\in\mathbb{R}^{3d\times d}$ is learnable parameter.
+5.2.3 Span Selection
+Upon acquiring the final representation $\tilde{\mathbf{h}}_{s{i}}$ for every slot across each event, we employ a methodology inspired by (Ma et al., 2022a) to convert these into a duo of span selectors $\Phi_{s_{i}}^{\mathrm{start}},\Phi_{s_{i}}^{\mathrm{end}}$:
+$\begin{array}[]{c}\Phi_{s_{i}}^{\mathrm{start}}=\tilde{\mathbf{h}}_{s{i}}\circ\mathbf{w}_{\mathrm{start}},\vspace{1.0ex}\ \Phi_{s_{i}}^{\mathrm{end}}=\tilde{\mathbf{h}}_{s{i}}\circ\mathbf{w}_{\mathrm{end}},\end{array}$ (14)
+where $\circ$ denotes element-wise multiplication and $\mathbf{w}_{\mathrm{start}},\mathbf{w}_{\mathrm{end}}\in\mathbb{R}^{d}$ are tunable parameters.
+These selectors, $\Phi_{s_{i}}^{\mathrm{start}}$ and $\Phi_{s_{i}}^{\mathrm{end}}$, are pivotal in identifying the start and end points of the slot $s_{i}$ within the original text.
+Following the approach in (Ma et al., 2022a), the Bipartite Matching Loss is utilized (Carion et al., 2020) to refine the alignment of predicted and actual argument spans during training ( For detailed content on this part, please refer to the work PAIE (Ma et al., 2022b), and we follow their approach here.)
 6 Experiments
@@
 6.1.1 Datasets
-We evaluate the proposed method on three large-scale document-level EAE datasets, that is, RAMS ebner2020multi , WikiEvents lietal2021document and MLEE 10.1093/bioinformatics/bts407 , and detailed data statistics are listed in Table 1.
+We evaluate the proposed method on three large-scale document-level EAE datasets, that is, RAMS (Ebner et al., 2020), WikiEvents (Li et al., 2021) and MLEE (Pyysalo et al., 2012), and detailed data statistics are listed in Table 1.
 Table 1: Detailed data statistics.
@@
     Test 871 365 2200
-  • 1.
   • RAMS  The dataset includes 9,124 annotated events sourced from English online news articles.
 In this dataset, each event forms a single instance, and events that occur within the same context are grouped together into a single instance containing multiple events.
-We adopt the original train/dev/test split following ebner2020multi .
-  • 2.
+We adopt the original train/dev/test split following (Ebner et al., 2020).
   • WikiEvents  This dataset centers on events documented in English Wikipedia, encompassing both the events and the related news articles referencing these events.
@@
 However, for our experiments, we only employ the precise argument annotations provided.
-  • 3.
   • MLEE  This dataset is a document-level event extraction dataset that focuses on bio-medical publications, consisting of manually annotated abstracts written in English.
-We follow the procedure outlined in 10.1093/bioinformatics/btaa540 and it’s worth noting that the preprocessed dataset only provides a train/test data split, without a dedicated development set.
-In this case, we use the training set as the development set for our experiments following  10.1093/bioinformatics/btaa540 .
+We follow the procedure outlined in (Trieu et al., 2020) and it’s worth noting that the preprocessed dataset only provides a train/test data split, without a dedicated development set.
+In this case, we use the training set as the development set for our experiments following  (Trieu et al., 2020).
 6.1.2 Evaluation Metrics
-For the RAMS dataset, we adopt the method detailed in xu2022two , focusing on both Span F1 and Head F1 scores across the development and test sets.
+For the RAMS dataset, we adopt the method detailed in (Xu et al., 2022b), focusing on both Span F1 and Head F1 scores across the development and test sets.
 Span F1 evaluates the precision of matching predicted argument spans against the gold standard, while Head F1 quantifies the accuracy of identifying the central word in each argument span.
-For a comprehensive assessment, we also incorporate Span F1 for argument identification (Arg IF) on the test set, as suggested by ma-etal-2022-prompt .
-Regarding the the WikiEvents dataset, we present both Head F1 and Coref F1 metrics on the test set for Argument Identification (Arg IF) and Argument Classification (Arg CF) tasks, consistent with the methodology described in lietal2021document .
-The Coref F1 score, inspired by the evaluation technique from jigrishman2008refining , gauges the coreference accuracy between extracted arguments and the established gold standard.
-For the MLEE dataset, we report Span F1 scores on the test set for both Arg IF and Arg CF tasks, following the methodologies in lietal2021document ; ma-etal-2022-prompt .
+For a comprehensive assessment, we also incorporate Span F1 for argument identification (Arg IF) on the test set, as suggested by (Ma et al., 2022b).
+Regarding the Wikievents and MLEE datasets, we report Span F1 scores on the test set for both Argument Identification (Arg IF) and Argument Classification (Arg CF) tasks, following the methodologies in (Li et al., 2021; Ma et al., 2022b).
 6.1.3 Baselines
 […]
-Table 2: Detailed hyperparameter settings.
+Table 2: Hyperparameter settings for $\text{CARLG}_{\text{span}}$.
+* means that we tuned the hyperparameters in our experiments.
+The rest of hyperparameters are set the same as TSAR (Xu et al., 2022b).
 Hyperparameters RAMS WikiEvents MLEE
@@
 Lambda 0.05 0.1 0.05
-Alpha 10 10 10
-Gamma 2 2 2
+Alpha* 10 10 10
+Gamma* 2 2 2
+Hyperparameters RAMS WikiEvents MLEE
+Training Steps 10000 10000 10000
+Warmup Ratio* 0.1 0.1 0.2
+Learning Rate* 2e-5 3e-5 3e-5
+Max Gradient Norm 5 5 5
+Batch Size* 4 4 4
+Context Window Size 250 250 250
+Max Span Length 10 10 10
+Max Encoder Seq Length 500 500 500
+Max Prompt Length* 210 360 360
+Encoder Layers* 17 17 17
+Decoder Layers* 7 7 7
+Table 3: Hyperparameter settings for $\text{CARLG}_{\text{prompt}}$.
+* means that we tuned the hyperparameters in our experiments.
+The rest of hyperparameters are set the same as PAIE (Ma et al., 2022a).
 6.1.4 Hyperparameter Setting
-Due to the fact that our proposed module does not rely heavily on hyperparameters, most of the hyperparameters follow the same configuration of TSAR xu2022two .
-More detailed hyperparameters settings are listed in Table 2.
+Due to the fact that our proposed module does not rely heavily on hyperparameters, most of the hyperparameters follow the same configuration of baselines.
+Here, we provide the hyperparameter settings for the span-based method TSAR in Table 2 and the prompt-based method PAIE in Table 3
 Table 4: The main results for the RAMS dataset.
 Arg IF denotes the F1 score for the Argument Identification subtask, and other metrics are for the Argument Classification subtask.
-* indicates results reported by  ma-etal-2022-prompt .
-The highest scores are highlighted in bold, while the second-highest scores are underlined.
+Bold indicates the best experimental results.
+The values in parentheses represent the performance compared to the original methods after incorporating our approach.
 Method Dev Test
 Span F1 Head F1 Span F1 Head F1 Arg IF
-Tag-based
-BERT-CRF ebner2020multi 38.1 45.7 39.3 47.1 -
+Generation-based Methods
+EEQA (Du & Cardie, 2020b) - - 46.7 - 48.7
+FEAE (Wei et al., 2021) - - 47.4 - -
+BART-Gen (Li et al., 2021) - - 48.6 57.3 51.2
+${\text{EA}^{2}\text{E}}$ (Zeng et al., 2022) 46.1 53.9 47.8 56.4 53.2
+EACE (Zhou et al., 2024) - - 50.1 56.3 -
+Span-based Methods
+Two-Step (Zhang et al., 2020d) 38.9 46.4 40.1 47.7 46.5
 […]
-Generation-based
-FEAE weietal2021trigger - - 47.40 - -
+TSAR (Xu et al., 2022b) 49.2 56.8 51.2 58.5 56.1
+SCPRG (Liu et al., 2023b) 50.2 57.2 51.9 59.1 -
+Prompt-based Methods
+PAIE (Ma et al., 2022b) 50.9 56.5 52.2 58.8 56.8
+RKDE (Hu et al., 2023) - - 50.3 55.1 -
+HRA (Ren et al., 2023) - - 48.4 - 54.6
+SPEAE (Nguyen, 2023) - - 52.5 - 57.0
+TabEAE (He et al., 2023) 51.9 57.8 51.8 59.0 56.7
+$\text{CARLG}_{\text{span}}$ (Ours)
+Two-Step+$\text{CARLG}_{\text{span}}$ 43.4 ($\uparrow$4.5) 50.6 ($\uparrow$4.2) 45.3 ($\uparrow$5.2) 52.8 ($\uparrow$5.1) 50.3 ($\uparrow$3.8)
+${\text{TSAR}}_{{\text{large}}}$+$\text{CARLG}_{\text{span}}$ 50.7 ($\uparrow$1.5) 58.1 ($\uparrow$1.3) 52.6 ($\uparrow$1.4) 59.9 ($\uparrow$1.4) 57.1 ($\uparrow$1.0)
+$\text{CARLG}_{\text{prompt}}$ (Ours)
+PAIE+$\text{CARLG}_{\text{prompt}}$ 52.8 ($\uparrow$1.9) 58.7 ($\uparrow$2.2) 54.4 ($\uparrow$2.2) 60.7 ($\uparrow$1.9) 59.1 ($\uparrow$2.3)
+TabEAE+$\text{CARLG}_{\text{prompt}}$ 53.1 ($\uparrow$1.2) 59.2 ($\uparrow$1.4) 54.3 ($\uparrow$2.5) 60.3 ($\uparrow$1.3) 58.6 ($\uparrow$1.9)
+6.2 Main Results
+In this section, we compare the effectiveness of our CARLG framework on three datasets. We validate the CARLG framework both on span-based and prompt-based methods and then conduct further analyses.
+6.2.1 Performance of Span-based Methods equipped with CARLG
+As shown in Table 4 and Table 5, both Two-Step and TSAR exhibit improved performance when equipped with $\text{CARLG}_{\text{span}}$.
+Two-Step yields an improvement of +5.2/+5.1 Span F1/Head F1 for Arg CF and +3.8 Span F1 for Arg IF on RAMS, +6.9/+6.1 Span F1 for Arg IF/Arg CF on WikiEvents, and +2.2/+1.9 Span F1 for Arg IF/Arg CF on MLEE.
+TSAR achieves notable improvements in Span F1/Head F1 for Arg CF on RAMS (+1.4/+1.4), Span F1 for Arg IF/Arg CF on WikiEvents (+1.5/+1.5), and Span F1 for Arg IF/Arg CF on MLEE (+1.3/+1.2).
+It is worth noting that the improvement of TSAR after incorporating CCA and RLIG modules is relatively modest.
+This could be attributed to the fact that the AMR graph already contains contextual semantic information, which may mitigate the need for additional contextual clue integration.
+These results illustrate that our two modules can be easily applied to other span-based methods with transformer encoders and provide a significant performance boost in document-level EAE tasks, which are compact and transplantable.
+6.2.2 Performance of Prompt-based Methods equipped with CARLG
+When equipped with $\text{CARLG}_{\text{prompt}}$, as illustrated in Table 4 and Table 5, the prompt-based methods PAIE and TabEAE show significant enhancements across all three datasets, surpassing previous generation-based and span-based models.
+Specifically, PAIE shows a boost of +1.1/+1.4 in Span F1/Head F1 for Arg CF and +2.3 in Span F1 for Arg IF on RAMS, +2.1/+1.7 in Span F1 for Arg IF/Arg CF on WikiEvents, and +2.0/+1.3 in Span F1 for Arg IF/Arg CF on MLEE.
+Meanwhile, TabEAE achieves remarkable improvement in Span F1/Head F1 for Arg CF on RAMS (+2.5/+1.3), in Span F1 for Arg IF/Arg CF on WikiEvents (+0.7/+1.1), and in Span F1 for Arg IF/Arg CF on MLEE (+1.1/+0.9).
+These results validate the effective adaptability of our CARLG framework to prompt-based EAE methods, showcasing enhanced extraction capabilities.
+This is attributed to its adept utilization of contextual clues and the integration of latent role representations enriched with semantic connections.
+Table 5: Main results of WikiEvents and MLEE datasets.
+Bold indicates the best experimental results.
+The values in parentheses represent the performance compared to the original methods after incorporating our approach.
+WikiEvents MLEE
+Method Arg IF Arg CF Arg IF Arg CF
+Span F1 Span F1 Span F1 Span F1
+Generation-based Methods
+EEQA (Du & Cardie, 2020b) 56.9 54.5 70.3 68.7
+${\text{EA}^{2}\text{E}}$ (Zeng et al., 2022) 68.8 63.7 70.6 69.1
+EACE (Zhou et al., 2024) 71.1 66.2 - -
+Span-based Methods
+Two-Step (Zhang et al., 2020d) 52.9 51.1 66.7 61.9
+[…]
+TSAR (Xu et al., 2022b) 70.8 65.5 72.3 71.3
+SCPRG (Liu et al., 2023b) 71.3 66.4 - -
+Prompt-based Methods
+PAIE (Ma et al., 2022b) 70.5 65.3 72.1 70.8
+RKDE (Hu et al., 2023) 69.1 63.8 - -
+HRA (Ren et al., 2023) 69.6 63.4 - -
+SPEAE (Nguyen, 2023) 71.9 66.1 - -
+TabEAE (He et al., 2023) 71.1 66.0 75.1 74.2
+$\text{CARLG}_{\text{span}}$ (Ours)
+Two-Step+$\text{CARLG}_{\text{span}}$ 59.8 ($\uparrow$6.9) 57.2 ($\uparrow$6.1) 68.9 ($\uparrow$2.2) 63.8 ($\uparrow$1.9)
+${\text{TSAR}}_{{\text{large}}}$+$\text{CARLG}_{\text{span}}$ 72.1 ($\uparrow$1.5) 66.8 ($\uparrow$1.5) 73.4 ($\uparrow$1.3) 72.0 ($\uparrow$1.2)
+$\text{CARLG}_{\text{prompt}}$ (Ours)
+PAIE+$\text{CARLG}_{\text{prompt}}$ 72.6 ($\uparrow$2.1) 67.0 ($\uparrow$1.7) 74.1 ($\uparrow$2.0) 73.1 ($\uparrow$1.3)
+TabEAE+$\text{CARLG}_{\text{prompt}}$ 72.1 ($\uparrow$1.0) 67.1 ($\uparrow$1.1) 76.2 ($\uparrow$1.1) 75.1 ($\uparrow$0.9)
+6.2.3 Further analyses
+Further analyzing the experimental results in Table 4 and Table 5, we observe that: (1) Our CARLG framework shows relatively larger improvement on simple models such as Two-step and PAIE.
+We hypothesize this may be because other components of complex models have synergistic or antagonistic effects on CCA and RLIG modules, constraining the enhancement potential of the CARLG modules and preventing them from fully unleashing their effectiveness.
+(2) The prompt-based methods yield the best results, with those equipped with the CARLG framework outperforming both generation-based and span-based approaches in terms of performance.
+[…]
+RAMS WikiEvents
+Method Params Arg CF Arg IF Arg IF Arg CF
+Span F1 Head F1 Span F1 Span F1 Span F1
+$\text{PAIE}_{\text{C}}$ 408.22M 52.81±0.43 58.74±0.36 57.03±0.87 72.62±0.98 67.02±0.85
+    -CCA 407.13M 51.28±0.60 56.67±0.51 55.43±0.48 70.16±0.87 65.87±1.06
+    -RLIG 406.14M 51.77±0.36 56.96±0.44 55.91±0.48 70.83±0.79 66.49±0.68
+    -both 405.12M 50.94±0.36 56.47±0.22 55.02±0.24 69.85±0.45 65.12±0.58
+[…]
+    -CCA 373.65M 40.24±0.84 47.66±0.67 46.98±0.26 55.48±1.87 53.65±1.02
+    -RLIG 371.68M 41.67±1.12 48.26±0.95 47.24±0.60 56.43±1.46 54.87±1.19
+    -both 370.63M 38.93±0.80 46.44±0.66 46.10±0.90 52.92±0.56 51.09±1.22
+6.3 Ablation Study
 […]
@@
 […]
-EACE zhou2024eace - - 50.1 56.3
-RKDE hu2023role - - 50.3 55.1
-Span-based
-Two-Step zhang2020two 38.9 46.4 40.1 47.7 -
-[…]
-${\text{TSAR}}_{{\text{base}}}$ xu2022two 45.23 51.70 48.06 55.04 -
-${\text{TSAR}}_{{\text{large}}}$ xu2022two 49.23 56.76 51.18 58.53 -
-${\text{SCPRG}}_{{\text{base}}}$ liu-etal-2023-enhancing-document 46.06 52.75 48.18 55.83 -
-${\text{SCPRG}}_{{\text{large}}}$ liu-etal-2023-enhancing-document 50.23 57.22 51.88 59.06 -
-Our Methods
-[…]
-${\text{CARLG}}_{{\text{base}}}$ 46.83 53.42 49.81 56.21 55.91
-${\text{CARLG}}_{{\text{large}}}$ 50.64 57.71 52.82 59.79 56.71
-Span-based+Ours
-Two-Step+Ours 43.42 50.58 45.27 52.79 50.34
-${\text{TSAR}}_{{\text{base}}}$+Ours 46.52 53.28 49.23 56.07 53.22
-${\text{TSAR}}_{{\text{large}}}$+Ours 50.61 57.96 52.52 59.77 57.03
-5.2 Main Results
-5.2.1 Performance of our enhanced baseline
-[…]
-5.2.2 Performance of CARLG
-[…]
-Table 4: Main results of WikiEvents and MLEE datasets.
-* means we rerun their code based on their experimental settings (hyperparameter tuning is conducted when necessary).
-The highest scores are highlighted in bold, while the second-highest scores are underlined.
-WikiEvents MLEE
-Method Arg IF Arg CF Arg IF Arg CF
-Head F1 Coref F1 Head F1 Coref F1 Span F1 Span F1
-Tag-based
-BERT-CRF ebner2020multi 69.83 72.24 54.48 56.72 - -
-Generation-based
-BERT-QA du-cardie-2020-event 61.05 64.59 56.16 59.36 - -
-BERT-QA-Doc du-cardie-2020-event 39.15 51.25 34.77 45.96 - -
-EEQA du-cardie-2020-event - - 56.9 - 68.4 66.7
-[…]
-[…]
-[…]
-${\text{EA}^{2}\text{E}}_{{\text{large}}}$ zengetal2022ea2e 74.62 75.77 68.61 69.70 - -
-EACE zhou2024eace - 71.1 - 66.2 - -
-RKDE hu2023role - 69.1 - 63.8 - -
-Span-based
-Two-Step zhang2020two 70.32* 72.27* 55.11* 57.12* 70.64* 68.74*
-${\text{TSAR}}_{{\text{base}}}$ xu2022two 75.52 73.17 68.11 66.31 72.21* 71.03*
-${\text{TSAR}}_{{\text{large}}}$ xu2022two 76.62 75.52 69.70 68.79 73.92* 72.71*
-${\text{SCPRG}}_{{\text{large}}}$ liu-etal-2023-enhancing-document 77.52 76.01 70.11 69.25 73.88* 73.21*
-Our Methods
-[…]
-${\text{CARLG}}_{{\text{base}}}$ 76.22 74.97 69.03 68.53 73.26 71.23
-${\text{CARLG}}_{{\text{large}}}$ 77.22 76.03 70.92 70.11 75.90 73.71
-Span-based+Ours
-Two-Step+Ours 73.20 72.96 60.90 61.04 72.42 70.11
-${\text{TSAR}}_{{\text{base}}}$+Ours 76.43 75.10 68.81 68.22 73.13 71.12
-${\text{TSAR}}_{{\text{large}}}$+Ours 77.46 76.10 70.88 70.28 74.98 73.48
-5.2.3 Performance of Span-based Methods equipped with CARLG
-In order to evaluate the compatibility of our CCA and RLIG modules, we equip these two modules on some classical span-based methods and conduct experiments on three datasets.
-As is shown in Table 3 and Table 4, both Two-Step and TSAR exhibit improved performance when equipped with CCA and RLIG modules.
-Two-Step yields an improvement of +5.17/+5.09 Span F1/Head F1 for Arg CF on RAMS, +2.88/+5.79 Head F1 for Arg IF/Arg CF on WikiEvents, and +1.78/+1.37 Span F1 for Arg IF/Arg CF on MLEE.
-${\text{TSAR}}_{{\text{large}}}$ achieves notable improvements in Span F1/Head F1 for Arg CF on RAMS (+1.34/+1.24), Head F1 for Arg IF/Arg CF on WikiEvents (+0.84/+1.18), and Span F1 for Arg IF/Arg CF on MLEE (+1.06/+0.77).
-It is worth noting that the improvement of TSAR after incorporating CCA and RLIG modules is relatively modest.
-This could be attributed to the fact that the AMR graph already contains contextual semantic information, which may mitigate the need for additional contextual clue integration.
-These results illustrate that our two modules can be easily applied to other span-based methods with transformer encoders and provide a significant performance boost in document-level EAE tasks, which are compact and transplantable.
-[…]
-Method Dev Test
-Span F1 Head F1 Span F1 Head F1
-[…]
- -Boundary Loss 42.18 50.28 42.33 51.68
- -Focal Loss 44.97 51.36 45.83 53.03
- -FIS 43.27 50.64 45.26 52.11
-Table 6: Ablation Study on RAMS and WikiEvents for CARLG. Span, Head and Coref refer to Span F1, Head F1 and Coref F1 scores respectively.
-RAMS WikiEvents
-Method Params Dev Test Arg IF Arg CF
-Span Head Span Head Head Coref Head Coref
-${\text{CARLG}}_{{\text{base}}}$ 120.87M 46.83 53.42 49.81 56.21 76.22 74.97 69.03 68.53
--CCA 120.28M 45.78 52.67 47.43 54.57 74.57 73.86 67.44 66.87
--RLIG 118.78M 45.27 52.36 47.91 54.52 75.59 73.83 68.26 67.49
--both 118.19M 45.07 51.59 45.76 53.16 73.92 73.22 66.90 65.98
-${\text{CARLG}}_{{\text{large}}}$ 374.70M 50.64 57.71 52.82 59.71 77.22 76.03 70.92 70.11
--CCA 373.65M 49.68 56.46 50.84 57.9 75.68 73.82 69.79 68.65
- -RLIG 371.68M 49.96 57.32 51.28 58.46 76.30 74.08 69.87 68.96
--both 370.63M 48.33 54.04 47.52 55.61 75.45 73.55 68.63 67.30
-5.3 Ablation Study
-In order to effectively showcase the effectiveness of our components, we carry out an ablation study on CARLG using the RAMS and WikiEvents datasets, as depicted in Table 6.
-First, when we remove the Contextual Clues Aggregation (CCA) module, we observe a decrease in Span F1 and Head F1 scores of ${\text{CARLG}}_{{\text{base}}}$/ ${\text{CARLG}}_{{\text{large}}}$ by 2.83/1.98 and 1.64/1.81 on the test set, respectively, which indicates that our CCA module plays a crucial role in capturing essential contextual clues for document-level EAE.
-Moreover, when removing the Role-based Latent Information Guidance (RLIG) module [^3] We additionally remove the respective role tokens that were included in the input sequence., we observe a significant decrease in performance for ${\text{CARLG}}_{{\text{base}}}$/ ${\text{CARLG}}_{{\text{large}}}$, with a drop of 1.9/1.54 in Span F1 and 1.69/1.07 in Head F1 on the RAMS test set.
-This demonstrates the effect of the RLIG module in capturing meaningful semantic correlations among roles, improving the performance of document-level EAE.
-Moreover, when both the CCA and RLIG modules are removed, the decline in performance is even more noticeable in comparison to excluding a single module.
-This illustrates the synergistic enhancement achieved by the collaboration of these two modules.
-[…]
-Figure 7: Context weights for a sample from RAMS are displayed, focusing on the span-trigger pair (Islamic State, wounded). Various color intensities are used to depict the attention weights.
+Figure 7: Attentive weights for an event instance from RAMS dataset.
+The event is life.die.death-caused-by-violent-events, focusing on the argument government and trigger bombarding).
+Various color intensities are used to depict the attention weights.
+We conduct visualization experiments on $\text{CARLG}_{\text{span}}$ and $\text{CARLG}_{\text{prompt}}$, selecting TSAR and PAIE as the backbone methods, respectively.
 6.4 Analysis of Contextual Attentive Weights
-To evaluate the validity of our Contextual Clues Aggregation (CCA) module in capturing beneficial contextual clues for candidate argument spans, we present a visualization of the contextual weights $\textbf{p}^{c}_{i:j}$, as detailed in Eq.12, using an instance from the RAMS dataset.
-As shown in Figure 7, the CCA module assigns high weights to non-argument contextual clues such as attack, responsibility, and terrorist attack, which are highly relevant to the span-trigger pair (Islamic State, wounded).
-Notably, our Contextual Clue Aggregation (CCA) module also grants relatively high attentive weights to some words in other arguments, including explosive, Dozens, and Kabul.
-This implies that these argument words offer crucial information for predicting the role of Islamic State.
-This visualization illustrates how our Contextual Clue Aggregation (CCA) module effectively identifies relevant contextual clues for candidate spans and models the interaction of information among interconnected arguments within an event.
+To evaluate the validity of our Contextual Clues Aggregation (CCA) module in capturing beneficial contextual clues for candidate argument spans, we present a visualization of the contextual weights $\textbf{p}^{C}_{a_{i}}$, as detailed in Eq.1, using an instance from the RAMS dataset.
+We conduct visualization experiments on $\text{CARLG}_{\text{span}}$ and $\text{CARLG}_{\text{prompt}}$, selecting TSAR and PAIE as the backbone methods, respectively.
+As shown in Figure 7, the CCA module assigns high weights to non-argument contextual clues such as surrender, starving, and shooting, which are highly relevant to the argument-trigger pair (government, bombarding).
+Notably, our Contextual Clue Aggregation (CCA) module also grants relatively high attentive weights to some words in other arguments, including shelling, people.
+This implies that these argument words offer crucial information for predicting the role of government.
+Additionally, we find that for $\text{CARLG}_{\text{prompt}}$, the CCA module also learns high attention towards prompt words such as killed and die, indicating that the CCA module assists the model in better understanding the prompt.
+This visualization illustrates how our CCA module effectively identifies relevant contextual clues for target arguments and models the interaction of information among interconnected arguments within an event.
 Figure 8: We provide visualizations of attention weights directed towards the context for various candidate spans within an event.
-Additionally, our investigation extends to examining the attention weights associated with various span-trigger pairings within a single event.
-In Figure 8, we display a heat map created from the attention weights of 30 randomly chosen candidate spans in an event, with respect to their contextual relevance.
+Additionally, our investigation extends to examining the attention weights associated with various argument-trigger pairings within a single event.
+In Figure 8, we display a heat map created from the attention weights of 30 randomly chosen candidate spans in an event, with respect to their contextual relevance [^5] We conduct this experiment on span-based EAE methods TSAR (Xu et al., 2022b) equipped with our $\text{CARLG}_{\text{span}}$.
 The heat map reveals that different candidate arguments pay attention to different contextual information, suggesting that our Contextual Clues Aggregation (CCA) module is capable of dynamically identifying and utilizing pertinent contextual clues specific to each candidate argument span.
 This adaptability highlights the module’s efficiency in tailoring its focus to suit the unique contextual requirements of each argument span within an event.
-Figure 9: The visualization of cosine similarity between role representations of two examples from RAMS in sub-figure (a) and MLEE dataset in sub-figure(b) respectively.
+(a) The similarity for $\text{CARLG}_{\text{span}}$ in RAMS
+(b) The similarity for $\text{CARLG}_{\text{prompt}}$ in RAMS
+(c) The similarity for $\text{CARLG}_{\text{span}}$ in MLEE
+(d) The similarity for $\text{CARLG}_{\text{prompt}}$ in MLEE
+Figure 9: The visualization of cosine similarity between role representations of two examples from RAMS and MLEE datasets, respectively.
 Figure 10: Here’s an instance of a t-SNE visualization sourced from the RAMS dataset.
+We conduct this experiment using PAIE equipped with our $\text{CARLG}_{\text{prompt}}$ framawork.
 The representations of roles and arguments originate from five distinct documents.
-In sub-figure (a), we utilize average pooling representations encoded by BERT, while in (b), we visualize the representations fused with latent role embeddings.
+In sub-figure (a), we utilize average pooling representations encoded by BART (Lewis et al., 2019), while in (b), we visualize the representations fused with latent role embeddings.
 6.5 Analysis of Role Correlations
-To verify the capability of our model in capturing semantic correlations between roles, we visualize the cosine similarity among latent role representations from events within the RAMS and MLEE datasets.
+To verify the capability of our model in capturing semantic correlations between roles, we visualize the cosine similarity among latent role representations.
+We select two events within the RAMS and MLEE datasets for $\text{CARLG}_{\text{span}}$ and $\text{CARLG}_{\text{prompt}}$ [^6] We select TSAR and PAIE as the backbone methods, respectively., respectively.
 Figure 9 presents the results, where roles such as origin and destination, attacker and target, ToLoc and FromLoc exhibit similar representations, indicating their semantic similarity.
 This visualization validates our model’s efficacy in capturing semantic correlations among roles.
-Additionally, to substantiate the effective guidance provided by role representations, we utilize t-SNE visualization JMLR:v9:vandermaaten08a to display arguments corresponding to two distinct roles that appear together in five diverse documents, alongside their respective latent role embeddings.
-Figure 10a illustrates that arguments associated with the same role but in distinct documents are dispersed throughout the embedding space.
+Additionally, to substantiate the effective guidance provided by role representations, we utilize t-SNE visualization (van der Maaten & Hinton, 2008) to display arguments corresponding to two distinct roles that appear together in five diverse documents, alongside their respective latent role embeddings.
+Figure 10a illustrates that arguments associated with the same role but in distinct documents are dispersed throughout the embedding space [^7] We conduct this experiment using PAIE (Ma et al., 2022b) equipped with our $\text{CARLG}_{\text{prompt}}$ framawork.
 This dispersion is attributed to variations in the target events and their contextual settings.
@@
 Our CARLG method combines computational efficiency with low complexity while both CCA and RLIG introduce few parameters.
-As shown in Table 6, for ${\text{CARLG}}_{{\text{large}}}$, CCA utilizes the pre-trained encoder’s attention heads and performs multiplication and normalization operations, adding only about 0.28% new parameters.
-RLIG introduces approximately 0.81% new parameters totally.
+As shown in Table 6, for CCA utilizes the attention heads of the pre-trained encoder and performs multiplication and normalization operations, adding less than 0.5% new parameters.
+RLIG introduces approximately 0.8% new parameters totally.
 This makes our parameter quantity approximately equivalent to the transformer-based encoder plus a MLP classification layer.
+Figure 11: The parameter size, training time, and inference time of TSAR and PAIE before and after incorporating CARLG.
 Table 7 reports the efficiency comparison of various span-based models on MLEE dataset.
-Our ${\text{CARLG}}_{{\text{base}}}$ is approximately 2 times faster than ${\text{TSAR}}_{{\text{base}}}$ and can save up to 30.9% of inference time with smaller model size.
-Additionally, compared to Two-Step, our CARLG only increases 15.9% inference time and 8% model parameters.
-Such high efficiency and performance makes CARLG a promising choice for practical applications of document-level EAE.
-Table 7: Time cost (in seconds) of different span-based models on MLEE dataset. Experiments are run for one epoch on one Tesla A100 GPU.
+After integrating our $\text{CARLG}_{\text{span}}$ framework, ${\text{TSAR}}_{\text{C}}$ shows an increase of approximately 14.2% in training time and 11.4% in inference time.
+With the addition of $\text{CARLG}_{\text{prompt}}$, ${\text{PAIE}}_{\text{C}}$ shows an increase of about 7.4% in both training and inference time.
+The increased training and inference time associated with $\text{CARLG}_{\text{span}}$ is attributed to the exhaustive nature of the span-based EAE approaches, necessitating the enumeration of all possible spans.
+Consequently, the operations of the CCA and RLIG modules escalate in frequency as the quantity of candidate spans expands.
+Table 7: The parameter size and time cost (in seconds) before and after integrating the PAIE and TSAR models with the CARLG framework on MLEE dataset.
+Experiments are run for one epoch on one Tesla A100 (40G) GPU.
 Model Params Training Time Inference Time
-Two-Step 111.96M 229.68 55.29
-${\text{TSAR}}_{{\text{base}}}$ 142.03M 465.84 92.79
-${\text{CARLG}}_{{\text{base}}}$ 120.87M 280.08 64.08
+TSAR 392.44M 465.84 72.79
+PAIE 405.12M 148.74 74.31
+$\text{TSAR}_{\text{C}}$ 395.23M ($\uparrow$0.71%) 532.08 ($\uparrow$14.22%) 81.08 ($\uparrow$11.39%)
+$\text{PAIE}_{\text{C}}$ 408.22M ($\uparrow$0.77%) 159.80 ($\uparrow$7.44%) 79.81 ($\uparrow$7.40%)
 6.7 Error Analysis
-Table 8: Error analysis on the RAMS test set. The best results are in bold font.
+Table 8: Error analysis on the RAMS test set.
+${\text{TSAR}}_{\text{C}}$ and ${\text{PAIE}}_{\text{C}}$ represent the original TSAR and PAIE methods equipped with our CARLG framework.
+The best results are in bold font.
 Model Wrong Span Over-extract Partial Overlap Wrong Role Total
 Two-Step 86 64 47 32 46 275
-${\text{TSAR}}_{{\text{base}}}$ 81 48 57 28 19 233
-${\text{CARLG}}_{{\text{base}}}$ 79 45 55 29 15 223
+TSAR 81 48 57 28 19 233
+${\text{TSAR}}_{\text{C}}$ (Ours) 79 45 55 29 15 223
+PAIE 61 46 59 34 25 225
+${\text{PAIE}}_{{\text{C}}}$ (Ours) 59 45 59 29 18 210
 Figure 12: An extraction case from RAMS test set base on ${\text{Roberta}}_{{\text{large}}}$. This case shows both CCA module and RLIG module contribute to the argument extraction.
 To conduct a more comprehensive examination of the errors made by various models and analyze the underlying reasons for these errors , we manually compared the predictions of 200 randomly selected examples from the RAMS test set against their golden annotations.
-During our analysis, we categorized the errors into five main categories following xu2022two .
+During our analysis, we categorized the errors into five main categories following (Xu et al., 2022b).
 Wrong Span refers to the case where a specific role is assigned to a span that does not overlap with the corresponding golden span.
-We observe that this type of errors often occur when negative words such as ”not” are involved or when there are coreference spans related to the golden span.
+We observe that this type of errors often occur when negative words such as “not” are involved or when there are coreference spans related to the golden span.
 The category labeled Over-extract pertains to instances where the model predicts an argument role that does not genuinely appear in the document.
 Some of the extracted spans are either sub-strings of the golden spans (Partial) or exhibit overlaps with them (Overlap).
-The ”Partial” error category is also commonly observed when there is punctuation, such as a comma, within the golden span.
+The “Partial" error category is also commonly observed when there is punctuation, such as a comma, within the golden span.
 Wrong Role refers to the model correctly identifies the golden span, but assigns the wrong argument role to it.
-We compare the errors of some span-based methods, that is, Two-Step and ${\text{TSAR}}_{{\text{base}}}$ and ${\text{CARLG}}_{{\text{base}}}$.
-As shown in Table 8, compared with Two-Step and ${\text{TSAR}}_{{\text{base}}}$, our CARLG model decreases the number of errors from 275 to 223 and effectively reduces Wrong Span errors, which can be attributed to the integration of contextual clues and latent role guidance.
+We compare the number of errors on TSAR and PAIE before and after adding the CARLG framework.
+As shown in Table 8, compared with Two-Step and TSAR, our ${\text{TSAR}}_{\text{C}}$ decreases the number of errors from 275 to 223 and effectively reduces Wrong Span errors, which can be attributed to the integration of contextual clues and latent role guidance.
+Meanwile, compared with the original PAIE method, the number of errors decreases from 225 to 210 after adding our $\text{CARLG}_{\text{prompt}}$, especially Wrong Role and Overlap errors.
+This fully demonstrates that our proposed RLIG module can capture semantic correlations between roles.
+Table 9: Comparison with large language model methods on the test set of RAMS.
 Method RAMS
@@
 gpt-4 50.4 42.8
-CARLG (Ours) 56.71 52.82
-Table 9: Comparison with large language model methods on the test set of RAMS.
+$\text{TSAR}_{\text{C}}$ (Ours) 57.1 52.6
+$\text{PAIE}_{\text{C}}$ (Ours) 59.1 54.4
 6.8 Comparison with Large Language Models
-Large language models (LLMs) have become a focal point of interest among researchers, given their remarkable versatility across a diverse range of Information Extraction (IE) tasks xu2023large ; li2023semi ; zhang2024ultra .
+Large language models (LLMs) have become a focal point of interest among researchers, given their remarkable versatility across a diverse range of Information Extraction (IE) tasks (Xu et al., 2023a; Li et al., 2023; Zhang et al., 2024).
 In this paper, we undertake a comprehensive comparison with some large language models on the document-level EAE task.
@@
 The experimental results are detailed in Table 9.
-We utilize the prompts defined in the work zhou2023heuristics and experiments are conducted utilizing three prominent large language models: text-davinci-003 ouyang2022training , gpt-3.5-turbo ouyang2022training and GPT-4 achiam2023gpt , all of which are accessed via OpenAI’s public APIs  [^4] https://openai.com/api/.
+We utilize the prompts defined in the work (Zhou et al., 2023) and experiments are conducted utilizing three prominent large language models: text-davinci-003 (Ouyang et al., 2022), gpt-3.5-turbo (Ouyang et al., 2022) and GPT-4 (Achiam et al., 2023), all of which are accessed via OpenAI’s public APIs  [^8] https://openai.com/api/.
 As illustrated in Table 9, LLMs exhibit a notable performance disparity compared to supervised learning models in the EAE task.
@@
 6.9 Case Study
-In order to explore how the CCA and RLIG modules work in actual cases, we conduct the case study for our CARLG.
+In order to explore how the CCA and RLIG modules work in actual cases, we conduct the case study for our CARLG [^9] We present a case study on $\text{CARLG}_{\text{span}}$, choosing TASR (Xu et al., 2022b) as the backbone model..
 As shown in Figure 12, equipped with CCA and RLIG modules, our CARLG can correctly predict all arguments, while TSAR fails to predict arguments attack.
@@
 7 Conclusion
-In this paper, we introduce the CARLG model for document-level event argument extraction (EAE).
+In this paper, we introduce the CARLG model for document-level event argument extraction.
 CARLG comprises two novel modules: Contextual Clues Aggregation (CCA) and Role-Based Latent Information Guidance (RLIG).
-CCA effectively aggregates non-argument clue words to capture beneficial contextual information, while RLIG provides guidance based on latent role representations that capture semantic correlations between some roles.
+CCA module effectively aggregates non-argument clue words to capture beneficial contextual information, while RLIG module provides guidance based on latent role representations that capture semantic correlations between some roles.
 Diverse experiments show that our CARLG model surpasses current SOTA EAE models on three large-scale document-level datasets.
-While our CARLG model has exhibited remarkable performance in document-level event argument extraction (EAE) tasks with known event triggers, we acknowledge its limitation in depending on the availability of event triggers.
+While our CARLG model has exhibited remarkable performance in document-level event argument extraction tasks with known event triggers, we acknowledge its limitation in depending on the availability of event triggers.
 In practical situations, event triggers may not always be supplied or known in advance.
@@
 Furthermore, we intend to explore the applicability of the CARLG framework in other information extraction tasks, such as relation extraction.
+[…]